### v4

# On $\mathcal{F}$-semi-transitive families of operators

Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [30], in this paper, we introduce the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples.
Keywords:
Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
1. Introduction
The research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [33, 35]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 23, 37]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [22, 27, 28]. Further, Furstenberg topological transitivity of weighted composition operators, as well as of some other classes of operators, has been studied in several papers, see [2, 7, 18, 19, 29]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [22]. These adjoints considered in [22] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. All these facts motivated us to study in this paper the Furstenberg-dynamics of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, we introduce first the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [30] and as an extension of the notion of $\mathcal{F}$--transitivity Then, in Theorem 3.1, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm. A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [22, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
As mentioned above, our concept of $\mathcal{F}$-semi-transitivity is motivated by the concept of topological transitivity for supercyclicity given in [30]. We recall in this context that supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators. For example, Hilden and Wallen in [20] proved that any unilateral backward weighted shift is supercyclic. Afterwards, Salas in [35] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion. Supercyclicity of several kinds of operators has also been studied in for instance [16, 30, 32, 38]. Moreover, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 9, 17]. On the other hand, after pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [24, 25, 39]. In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [25, 34], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [25, 34]. This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [25, 34]. In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [30], and then, as a continuation and extension of the results in [25, 34], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding Furstenberg family is finitely invariant. Once again it turns out that a special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.3 and Proposition 4.4 in this paper. At the end of the paper, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 4.5).
2. Preliminaries
Throughout this paper, we let $\Omega$ be a locally compact non-compact Hausdorff space and $\alpha$ be an a homeomorphism of $\Omega.$ As usual, $C_{0}(\Omega)$ denotes the space of all continuous functions on $\Omega$ vanishing at infinity, $C_{b}(\Omega)$ denotes the space of bounded, continuous functions on $\Omega$, whereas $C_{c}(\Omega)$ stands for the set of all continuous, compactly supported functions on $\Omega$. Both $C_{0}(\Omega)$ and $C_{b}(\Omega)$ are equipped with the supremum norm. Moreover, we let $w$ be a positive bounded function on $\Omega$ such that also $w^{-1}:=\frac{1}{w}\in C_{b}(\Omega)$ and we put then $T_{\alpha,w}$ to be the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega).$ Easily, one can see that by the above assumptions $T_{\alpha,w}$ is well-defined and $\|T_{\alpha,w}\|\leq\|w\|_{\sup}$. Since $\frac{1}{w}$ is also bounded, then $T_{\alpha,w}$ is invertible and we have
$T_{\alpha,w}^{-1}f=\frac{f\circ\alpha^{-1}}{w\circ\alpha^{-1}},\quad(f\in C_{0}(\Omega)).$
Simply we denote $S_{\alpha,w}:=T_{\alpha,w}^{-1}$.
By some calculation one can see that for each $n\in\mathbb{N}$ and $f\in C_{0}(\Omega)$,
$T^{n}_{\alpha,w}f=\left(\prod_{j=0}^{n-1}(w\circ\alpha^{j})\right)\cdot(f\circ\alpha^{n})\text{ (1)}$
and
$S^{n}_{\alpha,w}f=\left(\prod_{j=1}^{n}(w\circ\alpha^{-j})\right)^{-1}\cdot(f\circ\alpha^{-n}).\text{ (2)}$
The adjoint $T_{\alpha,w}^{\ast}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm. It is straightforward to check that
$T_{\alpha,w}^{\ast}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
${T_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
and
${S_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$
In this paper, $\mathcal{F}$ represents a Furstenberg family, which is a specific collection of subsets satisfying the following conditions:
  1. (1)
    The empty set $\emptyset$ is not an element of $\mathcal{F}$.
  2. (2)
    If $A$ belongs to $\mathcal{F}$ and $B$ contains $A$, then $B$ also belongs to $\mathcal{F}$.
3. $\mathcal{F}$-semi-transitive cosine operator functions
Assume that $S$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $S.$
Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$. For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$. We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C_{z}^{\ast}\ (U_{1})\cap U_{2}\neq\emptyset\}$
is an element of $\mathcal{F}$.
The next theorem gives some sufficient conditions for the family $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ to be $\mathcal{F}$-semi-transitive.
Theorem 3.1.
We have $(ii)\Rightarrow(i)$.
  1. $(i)$
    $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive;
  2. $(ii)$
    for each compact subset $K$ of $\Omega$, each $\varepsilon>0$, and any two non-zero measures $\mu$ $\nu\in M(\Omega)$ with $|\mu|(K^{c})=|\nu|(K^{c})=0$, there exists some $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ of $K$ such that
    $|\mu|(A_{z})<\varepsilon\text{, }|\nu|(A_{z})<\varepsilon\text{, }D_{z}\cap E_{z}=\emptyset\text{, }A^{c}_{z}\cap K=D_{z}\cup E_{z},$
    for all $z\in F$, and
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\sup_{t\in D_{z}}\left(\prod_{j=0}^{1}\left(w_{z}\circ\alpha_{z}^{j}\right)(t)\right)<\varepsilon,$
    $\sup_{t\in E_{z}}\left(\prod_{j=1}^{2}\left(w_{z}\circ\alpha_{z}^{-j}\right)^{-1}(t)\right)<\varepsilon\ \ \text{ for all }z\in F.$
Proof.
Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$. Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$. As in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$. Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$. Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
$\varepsilon:=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}.$
For each $z\in F$, let $\tilde{\mu_{z}}$, $\tilde{\nu_{z}}$, $\tilde{\tilde{\nu_{z}}}$ be the elements of $M(\Omega)$ given by $\tilde{\mu_{z}}(B)=\mu(B\cap A_{z}^{c}\cap K)$, $\tilde{\nu_{z}}(B)=\nu(B\cap D_{z})$, and $\tilde{\tilde{\nu_{z}}}(B)=\nu(B\cap E_{z})$ for every $z\in F$ and any Borel subset $B$ of $\Omega$. Then
$|\mu-\tilde{\mu_{z}}|(\Omega)=|\mu|(A_{z})<\varepsilon<\delta\ \ \text{ and }\ \ |\mu-\tilde{\nu_{z}}-\tilde{\tilde{\nu_{z}}}|(\Omega)=|\nu|(A_{z})<\varepsilon<\delta,$
[…]
$\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|>0\ \ \text{ and }\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|>0.$
For each $z\in F$, put
[…]
Then
$\|\phi_{z}-\mu\|\leq\|\mu-\tilde{\mu_{z}}\|+2\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}$
and
$\left\|\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})-\nu\right\|$ $\leq\frac{1}{2}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}\ \sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}$
$+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
for all $z\in F$. Now, by exactly the same arguments as in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
$\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\mu\|\right\}\text{,}$
$\left\|T^{\ast}_{z}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{w_{z}(t)\|\nu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\nu\|\right\}\text{,}$
$\left\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{\prod_{j=0}^{1}(w_{z}\circ\alpha_{z}^{j})(t)\|\nu\|\right\}\text{,}$
and
$\left\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{\prod_{j=1}^{2}(w_{z}\circ\alpha_{z}^{-j})^{-1}(t)\|\nu\|\right\}\text{.}$
Utilizing this fact together with assumptions $(ii)$ w. r. to $\varepsilon=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}$ and the fact that $B(\mu,\delta)\subset U_{1}$ and $B(\nu,\delta)\subset U_{2},$ we deduce that
$\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})\in U_{1}$
for all $z\in F$. Hence
$F\subseteq\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}.$
Since $\mathcal{F}$ is Furstenberg , we obtain that
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}\in\mathcal{F}.$
Hence, $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive since $U_{1}$ and $U_{2}$ were arbitrary. ∎
Corollary 3.2.
Let $w$ be a continuous positive function on $\Omega$ such that $w$ and $\frac{1}{w}$ are bounded and let $\alpha$ be a homeomorphism of $\Omega$. Suppose that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and either
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{2n-1}\left(w\circ\alpha^{j}\right)(t)\right)=0$
or
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0.$
Then, for each non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, there exists some $N\in\mathbb{N}$ and some sequence $\{\lambda_{n}\}_{n=N}^{\infty}$ of positive numbers such that $\lambda_{n}C_{n}^{\ast}(U_{1})\cap U_{2}\neq\emptyset\text{ for all }n\geq N,$ where $C_{n}^{\ast}=\frac{1}{2}\left(T^{\ast n}_{\alpha,w}+S^{\ast n}_{\alpha,w}\right)$.
Proof.
Let $S=\mathbb{N}$ and $\mathcal{F}$ be the family of all infinite subset of $\mathbb{N}$. For each $n\in\mathbb{N}$, we let $\alpha_{n}:=\alpha^{n}$ and $w_{n}:=\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)$. Then we consider the case when $A_{n}=E_{n}=\emptyset$ and $D_{n}=K$ (such that $\tilde{\tilde{\nu_{n}}}=0$) for all $n\in\mathbb{N}$, and the case when $A_{n}=D_{n}=\emptyset$ and $E_{n}=K$ for all $n\in\mathbb{N}$, (such that $\tilde{\nu_{n}}=0$ for all $n\in\mathbb{N}$). By adapting the proof of Theorem 3.1 for these special cases, we can deduce the statement of corollary. ∎
Now we will provide a concrete example where the conditions of Corollary 3.2 are satisfied.
Example 3.3.
Let $X=\mathbb{R}$ and $M,\delta$ be positive constants such that $M\geq 2+2\delta$ and $\delta\geq 1.$ Put $\alpha$ to be the function on $\mathbb{R}$ given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}.$ Set
$w(t)=\begin{cases}M\text{ for }t\leq-1,\\ M+\frac{t+1}{2}(1+\delta-M)\text{ for }t\in[-1,1],\\ 1+\delta\text{ for }t\geq 1.\\ \end{cases}.$
By some calculations one can check that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0,$
so the conditions of Corollary 3.2 are satisfied. However, it is not hard to check by some calculations that for every strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ and every $t\in\mathbb{R}$ it holds that
$\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
hence the conditions of [22, Proposition 3.3] are not satisfied. This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [22, Proposition 3.3].
4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
We start with the following definition.
Definition 4.1.
Let $S$ be a set and let $\mathcal{F}$ be a proper Furstenberg family of subsets of $S$ and $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, the set
$\{t\in S\ |\text{ there exists }\lambda_{t}\in\mathbb{R}^{+}\text{ such that }\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing\}$
belongs to $\mathcal{F}$.
In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$. We will denote by $\ell_{2}(\mathcal{C})$ the standard Hilbert module over $\mathcal{C}$, see [31, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\|x\|^{2}\sum_{j}y_{j}^{*}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
For each $m,J\in\mathbb{N}$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
[…]
As shown in the proof of [25, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. Further, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ let $\{W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(s)}$ has a bounded inverse. Also, assume that $\mathcal{U}^{(1)},\dots,\mathcal{U}^{(N)}$ are unitary operators on $H$. Then, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ we let $T_{n,s}$ be the operator on $\ell_{2}(\mathcal{\mathcal{C}})$ defined as
$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\}_{j\in\mathbb{Z}}.$
By [25], $T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{\mathcal{C}})$ for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\}$.
In what follows, we will assume that $\mathcal{F}$ is a Furstenberg family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
Now we are ready to give the main result of this section.
Theorem 4.2.
Under the above notation and assumptions, the following statements are equivalent.
(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\}.$
and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
Proof.
$(1)\Rightarrow(2)$ Let $J,m\in\mathbb{N}$ be given. Since the families $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-semi-transitive and $\mathcal{F}$ is finitely invariant. given $\varepsilon>0$, we can find some $F\subseteq\mathcal{F}$ with $\{1,\dots,2J\}\cap F=\varnothing$ such that for each $n\in F$, there exists some $x_{n}\in\ell_{2}(\mathcal{C})$ and some $\lambda_{n}\in\mathbb{R}^{+}$ with
$\|(x_{n})-\tilde{p}_{J,m}\|<\varepsilon,$
$\|\lambda_{n}(T_{n,\ell}x_{n})-\tilde{p}_{J,m}\|<\varepsilon\text{ for all }\ell\in\{1,\dots,N\}.$
This gives further that
$\|(x_{n})_{j}-P_{m}\|<\varepsilon,$
and
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n}\|<\varepsilon.$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$, because
$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n},$
and $j+r_{\ell}n>J$ for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $n\in F$.
Hence
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{(\ell)r_{\ell}n}\|<\varepsilon$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$. Further, since
$\|\lambda_{n}(T_{n,s}\,x_{n})_{j}-P_{m}\|<\varepsilon\quad\text{for all }j\in\mathbb{Z}\text{ with }-J\leq j\leq J\text{ and }s\in\{1,\dots,N\},$
it follows that
$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}-P_{m}\|<\varepsilon,$
however,
$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|$
$\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
once again because $j-r_{s}n<J-2J\leq-J.$ For each $n\in F,$ $\ell,s\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, define
$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}.$
Then
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\,\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
Finally, if $\ell,s\in\{1,\dots,N\}$ with $\ell\neq s$, then $|r_{\ell}-r_{s}|n>2J$ for all $n\in F$. Hence, either $j-r_{s}n+r_{\ell}n<-J$ or $j-r_{s}n+r_{\ell}n>J$ for any $j\in\mathbb{Z}$ with $J\leq j\leq J$. Therefore, $(\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}=0,$ hence
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\big\|$
$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(\ell)r_{\ell}n}\big\|$
$\big\|\lambda_{n}(T_{n,\ell}(x_{n}))_{j-r_{s}n+r_{\ell}n}\big\|=\big\|(\lambda_{n}T_{n,\ell}(x_{n})-\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}\big\|<\varepsilon.$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$, and $s,\ell\in\{1,\dots,N\}$ with $s\neq\ell$.
$(2)\Rightarrow(1)$ Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m}x=x$ and $\tilde{p}_{J,m}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}$ since the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For each $a\in\ell_{2}(\mathcal{C}),$ let $B_{\delta}(a)$ denote the open ball with centre in $a$ and radius $\delta$ . We can find some $\delta>0$ such that $B_{\delta}(x)\subseteq\mathcal{O}$ and $B_{\delta}(y^{(\ell)})\subseteq V_{\ell}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
$\varepsilon:=\frac{\delta}{2JN(\|x\|+2C_{y}+\delta)}.$
Choose $F\in\mathcal{F}$ and families
$\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$
for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
and
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-n}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}^{*(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
Since $\|D_{n,j}-P_{m}\|<\varepsilon$ and $\|G_{n,j}^{(\ell)}-P_{m}\|<\varepsilon<1$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, it follows that $D_{n,j}\neq 0$ and $G_{n,j}^{(\ell)}\neq 0,$ hence $W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\neq 0$, and $W_{n,j}^{(\ell)-1}G_{n,j}^{(\ell)}\neq 0$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$. Now we notice that
$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
because
$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
Moreover, since for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$,
$(T_{n,\ell}(v_{n}^{(s)}))_{j-r_{s}n+r_{\ell}n}=$
$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}^{*(s)\,r_{s}n}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
it follows that
$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
Finally, it is easy to see that by the construction of $v_{n}^{(\ell)}$, we have
$T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else.}$
For each $n\in F$, set
$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
$x_{n}\in\mathcal{O}\quad\text{and}\quad\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}T_{n,\ell}(x_{n})\in V_{\ell}$
for all $n\in F$ and $\ell\in\{1,\dots,N\}$.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$ and
$\{W_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{W_{j}^{(N)}\}_{j\in\mathbb{Z}}\subset B(H)$
be such that each $W_{j}^{(\ell)}$ $(j\in\mathbb{Z},\,1\leq\ell\leq N)$ is invertible, and such that
$\|W_{j}^{(\ell)}\|\leq M\quad\text{for all }j\in\mathbb{Z},\,\ell\in\{1,\dots,N\},$
and some fixed $M>0$. For each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$, we let $W_{n,i}^{(l)}\in B(H)$ be given by
$W_{n,i}^{(l)}=W_{i}^{(\ell)}W_{i-1}^{(\ell)}\cdots W_{i-r_{\ell}n+1}^{(\ell)}.$
Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [25]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}^{(\ell)}$, i.e.
[…]
Then it is not hard to check that for each $\ell\in\{1,\dots,N\}$ it holds that
[…]
We will now present an application of Theorem 4.2 in this concrete case.
Corollary 4.3.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators
$\{D_{j}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{j,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{j,N}^{(k)}\}_{k\in\mathbb{N}}.$
in $\mathcal{C}$ for every $j\in\mathbb{Z}$ with $J\leq j\leq J$ such that
$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
and
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0.$
for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|=0.$
In order to construct concrete examples satisfying the conditions $(2)$ of Corollary 4.3, we need first the following technical result.
Proposition 4.4.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0.,$
then he sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
The main idea in this proof is motivated by the proofs of [25, Theorem 3.2] and [26, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
$\{f_{i,j}^{(\eta)}\}_{i\in\mathbb{N}},\subseteq H_{0}^{(j)},\{h_{i,l,j}^{(\eta)}\}_{i\in\mathbb{N}}\subseteq H_{l}^{(j)}\text{ for each }\ell\in\{1,\dots,N\}$
such that
$f_{i,j}^{(\eta)}\to e_{\eta},\text{ and }h_{i,\ell,j}^{(\eta)}\to e_{\eta}\quad\text{as }i\to\infty,$
for all $\eta\in\{-m,\dots,m\}$ and $\ell\in\{1,\dots,N\}$. Since by the assumption we have for all $\ell,s\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ that
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}f_{i,j}^{(\eta)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ we have
$\lim_{k\to\infty}\big\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\dots W_{j-r_{\ell}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\dots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|=0,$
we can inductively easily construct a subsequence $\{n_{k_{i}}\}_{i\in\mathbb{N}}$ such that
$\big\|W_{j-r_{s}n_{k_{i}}}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k_{i}}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,f_{i,j}^{(\eta)}\big\|<\tfrac{1}{4m^{2}4^{i}},$
and
$\big\|W_{j+r_{\ell}n_{k_{i}}-r_{s}n_{k_{i}}}^{(\ell)}\cdots W_{j-r_{s}n_{k_{i}}+1}^{(\ell)}W_{j-r_{s}n_{k_{i}}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|<\tfrac{1}{2_{m}2^{i}}.$
For each $i\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$ we define operators $D_{j}^{(i)}$ and $G_{j,\ell}^{(i)}$ on $H$ by
$(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
By some calculations and arguments similar to those in the proof of [25, Theorem 3.2] and [26, Proposition 2.7], we can conclude that the sequences of operators
$\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
satisfy the conditions in (2) of Corollary 4.3. for every $j\in\mathbb{Z}$ with $J\leq j\leq J$. ∎
Now we are ready to give an example where the conditions of Corollary 4.3 and Proposition 4.4 are satisfied.
Example 4.5.
Let $N=1$ and $H=L^{2}(\mathbb{R})$. As in [25, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>0$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$. For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$. If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1$ for all $j<0$; or such that $w_{j}\chi_{[0,\infty)}\leq 1$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$, then by the same arguments as in [ ] one can show that the conditions of Proposition 4.4 are satisfied also when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R})$ and $j\in\mathbb{Z}$ with $H_{0}^{(j)}=H_{1}^{(j)}=\dots=H_{N}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z},$ where $C_{c}(\mathbb{R})$ stands for the space of all continuous compactly supported functions on $\mathbb{R}.$

### v5

@@
 Abstract.
-Motivated by the concept of disjoint topological transitivity for supercyclicity given in [30], in this paper, we introduce the concept of disjoint Furstenberg-semi-transitivity.
+Motivated by the concept of disjoint topological transitivity for supercyclicity given in [31], in this paper, we present the concept of disjoint Furstenberg-semi-transitivity.
 Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space.
-Also, we characterize disjoint Furstenberg-semi-transitive generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space.
+Also, we characterize disjoint Furstenberg-semi-transitive compositions of a right multiplier and a generalized bilateral weighted shift on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space.
 As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers.
@@
 1. Introduction
-The research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [33, 35].
-Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published.
-In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 23, 37].
-Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [22, 27, 28].
-Further, Furstenberg topological transitivity of weighted composition operators, as well as of some other classes of operators, has been studied in several papers, see [2, 7, 18, 19, 29].
-Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [22].
-These adjoints considered in [22] act on the space of Radon measures on a locally compact, non-compact Hausdorff space.
-All these facts motivated us to study in this paper the Furstenberg-dynamics of cosine operator functions generated by the adjoints of weighted composition operators.
-In Section 3 in this paper, we introduce first the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [30] and as an extension of the notion of $\mathcal{F}$--transitivity Then, in Theorem 3.1, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm.
-A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2.
-We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [22, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
-As mentioned above, our concept of $\mathcal{F}$-semi-transitivity is motivated by the concept of topological transitivity for supercyclicity given in [30].
-We recall in this context that supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators.
+Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators.
 For example, Hilden and Wallen in [21] proved that any unilateral backward weighted shift is supercyclic.
@@
 Moreover, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 9, 17].
-On the other hand, after pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [24, 25, 39].
-In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [25, 34], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [25, 34].
-This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [25, 34].
-In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [30], and then, as a continuation and extension of the results in [25, 34], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding Furstenberg family is finitely invariant.
-Once again it turns out that a special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.3 and Proposition 4.4 in this paper.
-At the end of the paper, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 4.5).
+On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 7, 19, 20, 30].
+Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [31] and characterize $\mathcal{F}$-semi-transitivity classes of operators on various Banach spaces.
+Now, the research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [35, 36].
+Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published.
+In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 24, 38].
+Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [23, 28, 29].
+Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [23].
+These adjoints considered in [23] act on the space of Radon measures on a locally compact, non-compact Hausdorff space.
+Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of these published papers.
+All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the adjoints of weighted composition operators.
+In Section 3 in this paper, after presenting the concept of $\mathcal{F}$-semi-transitivity, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm (Theorem 3.1).
+A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2.
+We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [23, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
+After pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [25, 26, 40].
+In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [26], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [26].
+This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [26].
+In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [31], and then, as a continuation and extension of the results in [26], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding $\mathcal{F}$-family is finitely invariant.
+Once again it turns out that a special case of this theory is the case when the $\mathcal{F}$-family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.4 and Proposition 4.6 in this paper.
+Also, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 4.8) where we also illustrate the difference between $\mathcal{F}$-semi-transitive (and thus also supercyclic) generalized weighted bilateral shift operators on Hilbert modules and supercyclic classical weighted bilateral shifts on the sequence space $\ell^{2}.$ In addition, we introduce a new concept of $\mathcal{F}$-C*-semi-transitivity for generalized bilateral shift operators on the standard Hilbert C*-module and we characterize disjoint $\mathcal{F}$-C*-semi-transitive generalized weighted bilateral shifts.
+By our very definition, $\mathcal{F}$-semi-transitivity implies $\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as illustrated in Example 4.9 in this paper.
+Next, hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 34].
+Motivated by those results, in Proposition 4.10 and Corollary 4.11 we provide sufficient conditions for disjoint $\mathcal{F}$-semi-transitivity of compositions of right multipliers and generalized weighted bilateral shift operators (where the corresponding weights are multiplied from the left) on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space and we illustrate these results with concrete example where these sufficient conditions are satisfied (Example 4.12).
+Finally, we conclude this paper with a remark (Remark 4.13) where we explain that our results from Section 4 remain valid also if we instead of a separable Hilbert space consider any standard Hilbert module over a unital C*-algebra.
 2. Preliminaries
@@
 Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$.
-As in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$.
+As in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$.
 Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$.
@@
 $+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
-for all $z\in F$. Now, by exactly the same arguments as in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
+for all $z\in F$. Now, by exactly the same arguments as in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
 $\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
@@
 $\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
-hence the conditions of [22, Proposition 3.3] are not satisfied.
-This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [22, Proposition 3.3].
+hence the conditions of [23, Proposition 3.3] are not satisfied.
+This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [23, Proposition 3.3].
 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
@@
 Definition 4.1.
-Let $S$ be a set and let $\mathcal{F}$ be a proper Furstenberg family of subsets of $S$ and $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$.
-We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, the set
-$\{t\in S\ |\text{ there exists }\lambda_{t}\in\mathbb{R}^{+}\text{ such that }\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing\}$
-belongs to $\mathcal{F}$.
+Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in S},\dots,\\ \{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$.
+We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
+$\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing.$
+We recall also that the families $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are said to be disjoint $\mathcal{F}$-transitive, or shortly $d\mathcal{F}$-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ we have
+$\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
 In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$.
 For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
-We will denote by $\ell_{2}(\mathcal{C})$ the standard Hilbert module over $\mathcal{C}$, see [31, Example 1.3.5].
+We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [32, Example 1.3.5].
 Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra.
@@
 $\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\|x\|^{2}\sum_{j}y_{j}^{*}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
-For each $m,J\in\mathbb{N}$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
+For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
 […]
-As shown in the proof of [25, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
+As shown in the proof of [26, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
 For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$.
 Further, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ let $\{W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(s)}$ has a bounded inverse.
-Also, assume that $\mathcal{U}^{(1)},\dots,\mathcal{U}^{(N)}$ are unitary operators on $H$.
+Also, assume that $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$.
 Then, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ we let $T_{n,s}$ be the operator on $\ell_{2}(\mathcal{\mathcal{C}})$ defined as
-$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\}_{j\in\mathbb{Z}}.$
+$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
 By [26], $T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{\mathcal{C}})$ for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\}$.
-In what follows, we will assume that $\mathcal{F}$ is a Furstenberg family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
-Now we are ready to give the main result of this section.
+In what follows, we will assume that $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
+Now we are ready to give the first result of this section.
 Theorem 4.2.
@@
 and
-$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\}.$
+$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\},$
 and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
@@
 and
-$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n}\|<\varepsilon.$
+$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon.$
 for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$, because
-$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n},$
+$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},$
 and $j+r_{\ell}n>J$ for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $n\in F$.
 Hence
-$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
-$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{(\ell)r_{\ell}n}\|<\varepsilon$
+$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
+$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon$
 for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$. Further, since
@@
 it follows that
-$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}-P_{m}\|<\varepsilon,$
+$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}-P_{m}\|<\varepsilon,$
 however,
-$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|$
-$\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
+$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|=\,\|W_{n,j}^{(s)-1}\frac{1}{\lambda_{n}}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|$
+$\|(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
 once again because $j-r_{s}n<J-2J\leq-J.$ For each $n\in F,$ $\ell,s\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, define
-$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}.$
+$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}.$
 Then
@@
 and
-$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\,\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
+$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|<\varepsilon,$
+$\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|=\|W_{n,j}^{(s)\,-1}\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|<\varepsilon,$
+so
+$\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
 Finally, if $\ell,s\in\{1,\dots,N\}$ with $\ell\neq s$, then $|r_{\ell}-r_{s}|n>2J$ for all $n\in F$.
@@
 Therefore, $(\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}=0,$ hence
-$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\big\|$
-$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(\ell)r_{\ell}n}\big\|$
+$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=$
+$\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\big\|$
+$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{\ell}^{r_{\ell}n}\big\|$
 $\big\|\lambda_{n}(T_{n,\ell}(x_{n}))_{j-r_{s}n+r_{\ell}n}\big\|=\big\|(\lambda_{n}T_{n,\ell}(x_{n})-\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}\big\|<\varepsilon.$
@@
 and
-$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-n}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}^{*(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
+$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
 Since $\|D_{n,j}-P_{m}\|<\varepsilon$ and $\|G_{n,j}^{(\ell)}-P_{m}\|<\varepsilon<1$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, it follows that $D_{n,j}\neq 0$ and $G_{n,j}^{(\ell)}\neq 0,$ hence $W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\neq 0$, and $W_{n,j}^{(\ell)-1}G_{n,j}^{(\ell)}\neq 0$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$.
@@
 because
-$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
+$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
 Moreover, since for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$,
 $(T_{n,\ell}(v_{n}^{(s)}))_{j-r_{s}n+r_{\ell}n}=$
-$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}^{*(s)\,r_{s}n}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
+$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}_{s}^{*\,r_{s}n}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
 it follows that
@@
 $x_{n}\in\mathcal{O}\quad\text{and}\quad\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}T_{n,\ell}(x_{n})\in V_{\ell}$
+for all $n\in F$ and $\ell\in\{1,\dots,N\}$. ∎
+Next, let $\{S_{n}\}_{n\in\mathbb{N}}$ be a sequence of bounded invertible linear operators on $H$ and $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be families of generalized weighted bilateral shifts on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
+For each $n\in F$ and $s\in\{1,\dots,N\}$ we let $S_{n}T_{n,s}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
+$\big(S_{n}T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big):=\{S_{n}W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
+It is not hard to check that the operator $S_{n}T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{C}).$ In fact, $S_{n}T_{n,s}$ is again a generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ with respective weights $\{S_{n}W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ for every $n\in F$ and $s\in\{1,\dots,N\}.$
+We will say that $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-C*-semi-transitive , or shortly $d\mathcal{F}$-C*-semi-transitive , if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $n\in F$ there exists some invertible bounded linear operator $S_{n}$ on $H$ satisfying that
+$\mathcal{O}\cap(S_{n}T_{n,1})^{-1}(V_{1})\cap\dots\cap(S_{n}T_{n,N})^{-1}(V_{N})\neq\varnothing.$
+The next theorem can easily be deduced from Theorem 4.2.
+Theorem 4.3.
+Under the above notation and assumptions, the following statements are equivalent.
+(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
+(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
+$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
+$\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|<\varepsilon,\text{ }\|W_{n,j}^{(s)\,-1}S_{n}^{-1}G_{n,j}^{(s)}\|<\varepsilon,$
+and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
+$\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
+Proof.
+For the proof of $(1)\Rightarrow(2),$ we can proceed in exactly the same way as in the proof of Theorem 4.2 part $(1)\Rightarrow(2)$ by considering $S_{n}$ and $S_{n}^{-1}$ instead of $\lambda_{n}$ and $\tfrac{1}{\lambda_{n}},$ respectively.
+For the proof of $(2)\Rightarrow(1),$ we can also proceed similarly as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$ However, the only difference is that in this case we let now for each $n\in F$, $v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
+$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else,}\end{cases}$
+which clearly gives that
+$\|(v_{n}^{(\ell)})\|\leq\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\|\|y^{(\ell)}\|.$
+Then it suffices to also notice in this case that for each $n\in F$ and $\ell\in\{1,\dots,N\}$ it holds that
+$\|S_{n}T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
+(where $u_{n}$ is the same as in the proof of Theorem 4.2 in the part $(2)\Rightarrow(1)$) and and moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and $n\in F$ we have
+$\|S_{n}T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
+Finally, by the construction of $v_{n}^{(\ell)}$, it follows that
+$S_{n}T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else,}$
+for each $n\in F$ and $\ell\in\{1,\dots,N\}.$
+For each $n\in F$, set
+$x_{n}:=u_{n}+\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
+By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
+$x_{n}\in\mathcal{O}\quad\text{and}\quad S_{n}T_{n,\ell}(x_{n})\in V_{\ell}$
 for all $n\in F$ and $\ell\in\{1,\dots,N\}.$
@@
 $W_{n,i}^{(l)}=W_{i}^{(\ell)}W_{i-1}^{(\ell)}\cdots W_{i-r_{\ell}n+1}^{(\ell)}.$
-Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [25]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}^{(\ell)}$, i.e.
+Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [26]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell}$, i.e.
 […]
@@
 […]
-We will now present an application of Theorem 4.2 in this concrete case.
+We will now present an application of Theorem 4.2 and Theorem 4.3 in this concrete case.
 Corollary 4.4.
-Under the above notation and assumptions, the following statements are equivalent:
+Under the above notation and assumptions, the following statements are equivalent.
 (1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators
-$\{D_{j}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{j,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{j,N}^{(k)}\}_{k\in\mathbb{N}}.$
-in $\mathcal{C}$ for every $j\in\mathbb{Z}$ with $J\leq j\leq J$ such that
+(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators in $\mathcal{C}$
+$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
+$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
+$\vdots$
+$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
+such that
 $\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
 and
-$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0.$
+$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0$
 for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
 $\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|=0.$
-In order to construct concrete examples satisfying the conditions $(2)$ of Corollary 4.3, we need first the following technical result.
+Corollary 4.5.
+Under the above notation and assumptions, the following statements are equivalent.
+(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
+(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N},$ a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ and sequences of operators in $\mathcal{C}$
+$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
+$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
+$\vdots$
+$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
+such that
+$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
+and
+$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0,$
+$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0,$
+for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$ and satisfying in addition that for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
+$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0.$
+In order to construct concrete examples satisfying the conditions in $(2)$ of Corollary 4.4 and Corollary 4.5, we need first the following technical results.
 Proposition 4.6.
@@
 and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
-$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0.,$
-then he sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
+$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0,$
+then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
 Proof.
-The main idea in this proof is motivated by the proofs of [25, Theorem 3.2] and [26, Proposition 2.7].
+The main idea in this proof is motivated by the proofs of [26, Theorem 3.2] and [27, Proposition 2.7].
 Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
 $(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
-By some calculations and arguments similar to those in the proof of [25, Theorem 3.2] and [26, Proposition 2.7], we can conclude that the sequences of operators
+By some calculations and arguments similar to those in the proof of [26, Theorem 3.2] and [27, Proposition 2.7], we can conclude that the sequences of operators
 $\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
 satisfy the conditions in (2) of Corollary 4.4. for every $j\in\mathbb{Z}$ with $J\leq j\leq J$. ∎
-Now we are ready to give an example where the conditions of Corollary 4.3 and Proposition 4.4 are satisfied.
+Similarly, based on Theorem 4.3 and Corollary 4.5, we can prove the following proposition.
+Proposition 4.7.
+Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H,$ a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
+$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}\,y_{s}^{(j)}\big\|=0,$
+$\lim_{k\to\infty}\big\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
+and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
+$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}y_{s}^{(j)}\|=0,$
+then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
+Thanks to Proposition 4.6 we will now give an example where we will show that it can happen that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}$ and that at a same time the sequence of operators $\{\tilde{T}^{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$-semi-transitive, which illustrates the important difference between $\mathcal{F}$-semi-transitivity (and thus supercyclicity) of generalized bilateral weighted shifts on Hilbert C*-modules and supercyclicity of the classical bilateral weighted shifts on the sequence space $\ell^{2}.$
+Example 4.8.
+Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$.
+As in [26, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
+Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$.
+For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$.
+If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [26, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
+By our definition, $d\mathcal{F}$-semi-transitivity implies $d\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as the next example illustrates.
+Example 4.9.
+Let again $N=1,r_{1}=1,H=L^{2}(\mathbb{R})$ and $\alpha$ be the function on $\mathbb{R}$ given by $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ Let $v=2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}$ and $\tilde{S}$ be the weighted composition operator on $H$ with respect to $v$ and $\alpha,$ that is $\tilde{S}(f)=v\cdot(f\circ\alpha)$ for all $f\in H.$ For each $n\in\mathbb{N},$ put $S_{n}:=\tilde{S}^{n}.$ If we let for each $j\in\mathbb{Z},$ $W_{j}$ be the identity operator on $H,$ then it is not hard to check that the conditions of Proposition 4.7 are satisfied with $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ Thus, the generalized non-weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ is $\mathcal{F}$-C*-semi-transitive, however, by Theorem 4.2 it can not be $\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
+Hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 34].
+Motivated by those results, in the rest of this paper, we shall study the dynamics of a composition of a right multiplier and generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ defined in the beginning of this section.
+To this end, let $\tilde{W_{1}},\dots,\tilde{W_{N}}$ be invertible, bounded linear operators on $H.$ For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $R_{s,n}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
+$R_{s,n}(\{x_{i}\}_{i\in\mathbb{Z}})=(\{x_{i}\cdot\mathcal{U}_{s}^{*\,r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{i\in\mathbb{Z}})$
+for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [32, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
+For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
+$\big((T_{n,s}\circ R_{s,n})(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
+We will now present an extension of Theorem 4.2 in this setting.
+Proposition 4.10.
+Under the above notation and assumptions, we have $(2)\Rightarrow(1).$
+(1) The families of operators $\{T_{n,1}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
+(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that we can find families
+$\{\tilde{D_{n}}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},$
+$\{D_{n,-J}\}_{n\in F},\{G_{n,-J}^{(1)}\}_{n\in F},\dots,\{G_{n,-J}^{(N)}\}_{n\in F},$
+$\{D_{n,-J+1}\}_{n\in F},\{G_{n,-J+1}^{(1)}\}_{n\in F},\dots,\{G_{n,-J+1}^{(N)}\}_{n\in F},$
+$\vdots$
+$\{D_{n,J}\}_{n\in F},\{G_{n,J}^{(1)}\}_{n\in F},\dots,\{G_{n,J}^{(N)}\}_{n\in F},$
+of compact operators on $H$ satisfying
+$\|\tilde{D_{n}}-P_{m}\|<\varepsilon,\|D_{n,j}-P_{m}\|<\varepsilon,\|\tilde{G_{n}}^{(s)}-P_{m}\|<\varepsilon,\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
+and
+$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\|<\varepsilon^{2}$
+for all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\};$ and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ it holds that
+$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|<\varepsilon.$
+Proof.
+Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$.
+Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [32, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
+$\varepsilon:=\frac{\delta}{(2JN+3)(\|x\|+2C_{y}+\delta)}.$
+Choose $F\in\mathcal{F}$ and families
+$\{\tilde{D_{n}}\}_{n\in F},\{D_{n,j}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},\{G_{n,j}^{(N)}\}_{n\in F}$
+for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
+For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
+$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j}\tilde{D_{n}},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
+and
+$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
+By similar arguments as in the proof of Theorem 4.2 part $(2)\Rightarrow(1)$, we can deduce that $\tilde{D_{n}}\neq 0$ and $\tilde{G_{n}}^{(s)}\neq 0$ for all $n\in F,s\in\{1,\dots,N\},$ and moreover
+$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|x\|$
+and
+$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|y^{(s)}\|.$
+For each $n\in F$, set
+$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|\|\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
+and proceed further as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$
+Let again, for each $\ell\in\{1,\dots,N\},$ $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell},$ as defined in the beginning of this section.
+Recalling that the involution on $\mathcal{C}$ is an isometry, by similar arguments as in the proof of Proposition 4.6, we easily obtain the following corollary of Proposition 4.10 in the special case when $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$
+Corollary 4.11.
+Under the above notation and assumptions, if there exist dense subsets
+$\tilde{H_{0}},\tilde{H_{1}},\dots,\tilde{H_{N}}$,
+$H_{0}^{(-J)},H_{1}^{(-J)},\dots,H_{N}^{(-J)},$
+$\vdots$
+$H_{0}^{(J)},H_{1}^{(J)},\dots,H_{N}^{(J)}$
+of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ each $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $\tilde{x}\in\tilde{H_{0}},x^{(j)}\in H_{0}^{(j)}$, $\tilde{y_{s}}\in\tilde{H_{s}},y_{s}^{(j)}\in H_{s}^{(j)}$ we have
+$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\|((\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}\tilde{x}\|=0,$
+and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
+$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}(\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|=0,$
+then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
+We end this paper with an example where the conditions of Corollary 4.11 are satisfied.
 Example 4.12.
 Let $N=1$ and $H=L^{2}(\mathbb{R})$.
-As in [25, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>0$.
-Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$.
-For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$.
-If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1$ for all $j<0$; or such that $w_{j}\chi_{[0,\infty)}\leq 1$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$, then by the same arguments as in [ ] one can show that the conditions of Proposition 4.4 are satisfied also when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R})$ and $j\in\mathbb{Z}$ with $H_{0}^{(j)}=H_{1}^{(j)}=\dots=H_{N}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z},$ where $C_{c}(\mathbb{R})$ stands for the space of all continuous compactly supported functions on $\mathbb{R}.$
+For each $j\in\mathbb{Z},$ let $W_{j}$ be the identity operator on $H.$ Put $\tilde{\alpha}$ to be the function on $\mathbb{R}$ given by $\tilde{\alpha}(t)=t-1$ for all $t\in\mathbb{R}$ and $\tilde{w}==2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}.$ If we let $\tilde{W}$ be the weighted composition operator on $H$ with respect to $\tilde{\alpha}$ and $\tilde{w}$, then by some calculations it can be checked that the the conditions of Corollary 4.11 are satisfied with $\tilde{H_{0}}=\tilde{H_{1}}=H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$
+Remark 4.13.
+We notice that due to [32, Proposition 2.2.1] our results in this section remain valid if we instead of a separable Hilbert space $H$ consider any standard Hilbert module $\ell_{2}(\mathcal{A})$ over a unital C*-algebra $\mathcal{A}.$
+Acknowledgement
+I am grateful to Professor Dragan S.
+Đorđević for suggesting to me Hilbert $C^{*}$-modules as the main field of my research and for introducing to me the relevant literature.
+Also, I am grateful to Professor S.M.
+Tabatabaie for suggesting to me linear dynamics of operators as the topic of my research and for introducing to me the relevant literature.

### v6

@@
 Finally, we illustrate our results with concrete examples.
+We end this paper with an algebraic generalization to the case of non-unital Banach algebras.
 Keywords:
@@
 A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2.
-We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [23, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
+We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [24, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
 After pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [26, 27, 41].
@@
 Finally, we conclude this paper with a remark (Remark 4.13) where we explain that our results from Section 4 remain valid also if we instead of a separable Hilbert space consider any standard Hilbert module over a unital C*-algebra.
+Section 5 provides algebraic generalization of Theorem 4.2 in this paper to the case of more general non-unital Banach algebras.
 2. Preliminaries
@@
 Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$.
-As in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$.
+As in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$.
 Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$.
@@
 $+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
-for all $z\in F$. Now, by exactly the same arguments as in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
+for all $z\in F$. Now, by exactly the same arguments as in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
 $\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
@@
 $\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
-hence the conditions of [23, Proposition 3.3] are not satisfied.
-This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [23, Proposition 3.3].
+hence the conditions of [24, Proposition 3.3] are not satisfied.
+This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [24, Proposition 3.3].
 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
@@
 Definition 4.1.
-Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in S},\dots,\\ \{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$.
-We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
+Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\\ \{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$.
+We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
 $\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing.$
@@
 For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
-We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [32, Example 1.3.5].
+We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [33, Example 1.3.5].
 Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra.
@@
 […]
-As shown in the proof of [26, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
+As shown in the proof of [27, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
 For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$.
@@
 (1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
+(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in F}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
 $\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
@@
 Proof.
-The main idea in this proof is motivated by the proofs of [26, Theorem 3.2] and [27, Proposition 2.7].
+The main idea in this proof is motivated by the proofs of [27, Theorem 3.2] and [28, Proposition 2.7].
 Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
 $(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
-By some calculations and arguments similar to those in the proof of [26, Theorem 3.2] and [27, Proposition 2.7], we can conclude that the sequences of operators
+By some calculations and arguments similar to those in the proof of [27, Theorem 3.2] and [28, Proposition 2.7], we can conclude that the sequences of operators
 $\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
@@
 Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$.
-As in [26, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
+As in [27, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
 Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$.
 For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$.
-If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [26, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
+If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [27, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
 By our definition, $d\mathcal{F}$-semi-transitivity implies $d\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as the next example illustrates.
@@
 $R_{s,n}(\{x_{i}\}_{i\in\mathbb{Z}})=(\{x_{i}\cdot\mathcal{U}_{s}^{*\,r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{i\in\mathbb{Z}})$
-for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [32, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
+for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [33, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
 For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
@@
 Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$.
-Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [32, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
+Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [33, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
 $\varepsilon:=\frac{\delta}{(2JN+3)(\|x\|+2C_{y}+\delta)}.$
@@
 Remark 4.13.
-We notice that due to [32, Proposition 2.2.1] our results in this section remain valid if we instead of a separable Hilbert space $H$ consider any standard Hilbert module $\ell_{2}(\mathcal{A})$ over a unital C*-algebra $\mathcal{A}.$
-Acknowledgement
-I am grateful to Professor Dragan S.
-Đorđević for suggesting to me Hilbert $C^{*}$-modules as the main field of my research and for introducing to me the relevant literature.
-Also, I am grateful to Professor S.M.
-Tabatabaie for suggesting to me linear dynamics of operators as the topic of my research and for introducing to me the relevant literature.
+We notice that due to [33, Proposition 2.2.1] our results in this section remain valid if we instead of a separable Hilbert space $H$ consider any standard Hilbert module $\ell_{2}(\mathcal{A})$ over a unital C*-algebra $\mathcal{A}.$
+5. Algebraic generalizations and axiomatic approach
+We recall first the following definition.
+Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$.
+We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$.
+We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
+$\|ba\|\leq\|b\|_{1}\|a\|$
+for all $a\in\mathcal{A}$ and $b\in\mathcal{A}_{1},$ where  $\|\text{ }\|_{1}$ denotes the norm on $\mathcal{A}_{1}.$
+Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}$ satisfying that given any open subset $O$ of $\mathcal{A}$ and $x\in O$ there exists some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}^{3}x\in O$ as well.
+This condition will be called the condition $(\mathcal{P}.)$
+For a set $\mathcal{S},$ let $\mathcal{F}$ be a family of subsets of $\mathcal{S}$, and let
+$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
+be families of isometric algebra isomorphism of $\mathcal{A}_{1}$ such that
+$\Phi_{t,j}(\mathcal{A})=\mathcal{A}\qquad(t\in\mathcal{S},\ 1\leq j\leq N).$
+and such that $\Phi_{t,j|\mathcal{A}}$ is an isometry for all $j\in\{1,\dots,N\}.$ We will assume that the system
+$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
+is disjoint aperiodic with respect to $\{p_{\alpha}\}_{\alpha}$, that is for each fixed $p_{\alpha}$ and every $H\in\mathcal{F}$ there exist some $F\subseteq H$ with $F\in\mathcal{F}$ such that
+$p_{\alpha}\,\Phi_{t,\ell}(p_{\alpha})=0,\quad p_{\alpha}\,\Phi_{t,\ell}^{-1}(p_{\alpha})=0,\quad\text{and}\quad p_{\alpha}\,\Phi_{t,\ell}\!\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)=0\quad\text{for all }t\in F$
+$\text{ and all }r,\ell\in\{1,\dots,N\}\text{ with }r\neq\ell.$ Further, we will assume that for each $\alpha$ and all $t\in\mathcal{S}$, $r,\ell\in\{1,\dots,N\}$, and $a\in\mathcal{A}$, it holds that
+$\|a\,p_{\alpha}\|\leq\|a\|,\quad\|a\,\Phi_{t,\ell}(p_{\alpha})\|\leq\|a\|,$
+$\|a\,\Phi_{t,\ell}^{-1}(p_{\alpha})\|\leq\|a\|,\quad\quad\|a\,\Phi_{t,\ell}\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)\|\leq\|a\|.$
+This condition will be called the condition (R) throughout this section.
+We now set up families $\{b_{t,1}\}_{t\in\mathcal{S}},\cdots,\{b_{t,N}\}_{t\in\mathcal{S}}$ of elements in $\mathcal{A}_{1}$, and for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$, we define the operator
+$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Phi_{t,\ell}(a).$
+Due to the condition $(E)$, it is not hard to check that $T_{t,\ell}$ is a bounded linear operator on $\mathcal{A}$.
+Further, we will assume that for each $\alpha$ there exists so called $\alpha$-inverse” for $b_{t,\ell}$ denoted by $b_{t,\ell,\alpha}^{-1},$ that is for each $\alpha$ there exists an element $b_{t,\ell,\alpha}^{-1}\in\mathcal{A}_{1}$ satisfying that $b_{t,\ell}b_{t,\ell,\alpha}^{-1}ap_{\alpha}=b_{t,\ell,\alpha}^{-1}b_{t,\ell}ap_{\alpha}=ap_{\alpha}$ for all $a\in\mathcal{A}.$
+Also, we will assume that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0.$ Such condition will be called the condition (C) throughout the manuscript.
+5.1. Main results
+We have the following theorem.
+Theorem 5.1.
+Under the above notation and assumptions, the following statements are equivalent.
+  1. (1)
+    The families $\{T_{t,1}\}_{t\in\mathcal{S}},\cdots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are $d\mathcal{F}$-semi-transitive.
+  2. (2)
+    For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t,1}\}_{t\in F}\dots,\{g_{t,N}\}_{t\in F}$ in $\mathcal{A}$ such that $\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\quad\|g_{t,l}-p_{\alpha}^{2}\|<\varepsilon,$ and
+    $\|b_{t,r,\alpha}^{-1}g_{t,r}\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}\quad\text{for all }t\in F\text{ and }r,l\in\{1,\ldots,N\}.$
+    Moreover, for each distinct $r,l\in\{1,\ldots,N\}$ and $t\in F$ it holds that
+    $\|\Phi_{t,r}^{-1}(b_{t,r})\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1}g_{t,l})\|<\varepsilon.$
+Proof.
+We prove first (1) $\Rightarrow$ (2).
+Since the families $\{T_{\Phi_{t,1},b_{t,1}}\}_{t\in S},\dots,\{T_{\Phi_{t,N},b_{t,N}}\}_{t\in S}$ are $d\mathcal{F}$-semi-transitive, given $\varepsilon>0$ and some fixed $p_{\alpha}$ we can find $a_{t}\in\mathcal{A}$ and $\lambda_{t}\in\mathbb{R}^{+}$ such that $\|a_{t}-p_{\alpha}\|<\varepsilon$ and $\|\lambda_{t}T_{\Phi_{t,l}},b_{t,l}(a_{t})-p_{\alpha}\|<\varepsilon\text{ for all }t\in H\text{ and some }H\in\mathcal{F}.$ This gives $\bigl\|\,(a_{t}-p_{\alpha})\,p_{\alpha}\,\bigr\|\;\leq\;\bigl\|a_{t}-p_{\alpha}\bigr\|\;<\;\varepsilon.$
+Moreover, since $\bigl\{\ {\Phi_{t,l}}\bigr\}_{t\in S,\;1\leq l\leq N}$ is a disjoint aperiodic system, we can find some $F\in\mathcal{F}$ with $F\subseteq H$ such that $p_{\alpha}\!\Phi_{t,l}^{-1}(p_{\alpha})=0,\;p_{\alpha}\!\Phi_{t,l}(p_{\alpha})=0,\;$ and $\Phi_{t,r}(\Phi_{t,l}^{-1}(p_{\alpha}))=0,\;\text{for all }\;t\in F\;\text{ and }r,l\in\{1,\dots N\}\text{ with }r\neq l.$ Hence, we get
+$\qquad\frac{1}{|\lambda_{t}|}\,\left\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\,\lambda_{t}\right\|=\left\|\!b_{t,r,\alpha}^{-1}\,\!\bigl(b_{t,r}\,\Phi_{t,r}(a_{t})\,(p_{\alpha})\,\bigr)\lambda_{t}\right\|$
+$\left\|a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|=\left\|(a_{t}-p_{\alpha})\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|\leq\left\|a_{t}-p_{\alpha}\right\|<\varepsilon.$
+and
+$\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,p_{\alpha}\;-\;p_{\alpha}^{2}\right\|\leq\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})-p_{\alpha}\right\|=\left\|\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha}\right\|<\epsilon,\qquad$
+for all $t\in F$ and $r\in\{1,\dots,N\}$.
+For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$
+Then $\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
+$\text{Similarly, for each }t\in F\text{ and }l\in\{1,\dots,N\},\text{ we have}$
+$\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})\,\Phi_{t,l}(p_{\alpha})\|\;=\;\|(\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})-p_{\alpha})\,\Phi_{t,l}(p_{\alpha})\|\;\leq\;\|\lambda_{t}\,T_{t,l}(a_{t})-p_{\alpha}\|\;<\;\varepsilon,$
+[…]
+$\text{ set }d_{t}=a_{t}p_{\alpha}.$
+$\text{Then }\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
+Hence, for all $t\in F$ and $r,l\in\{1,\dots,N\},$ we have
+$\|b_{t,r,\alpha}^{-1}(g_{t,r})\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}.$
+Further, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we get
+$\Bigl\|\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}\bigr)\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\Bigr\|\;=\;\Bigl\|\,\!b_{t,r}\,\Phi_{t,r}\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}g_{t,l}\bigr)\Bigr\|$
+$\bigl\|\,\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|$
+$\bigl\|\,(\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|\;<\;\varepsilon,$
+where we have used that $p_{\alpha}\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)=0\quad\text{for all }t\in F\text{ and each distinct }r,l\in\{1,\dots,N\}.$
+Now we prove the implication $(2)\Rightarrow(1)$.
+Let $O,V_{1},\dots,V_{N}$ be non-empty open subsets of $\mathcal{A}$.
+Then we can find some non-zero $x\in O$ and non-zero $y_{l}\in V_{l}$ for each $l\in\{1,\dots,N\}$.
+There exists some $\alpha$ such that $p_{\alpha}^{3}x\in O$ and $p_{\alpha}^{3}y_{l}\in V_{l}\setminus\{0\}$ for each $l=1,\dots,N$.
+Clearly, this gives that $p_{\alpha}x\neq 0$ and $p_{\alpha}y_{l}\neq 0$ for each $l=1,\dots,N$.
+Also, since $O\setminus\{0\},\,V_{1}\setminus\{0\},\dots,V_{N}\setminus\{0\}$ are open and non-empty, there is some $\delta>0$ such that the $\delta$-neighbourhood of $p_{\alpha}^{3}x$ is contained in $O\setminus\{0\}$, and the $\delta$-neighbourhood of $p_{\alpha}^{3}y_{l}$ is contained in $V_{l}\setminus\{0\}$ for each $l=1,\dots,N$.
+Let $C_{y}=\max_{1\leq l\leq N}\bigl\{\|y_{l}\|,\bigr\}$ and set
+$\varepsilon:=\min\!\left\{\frac{\delta}{4\|x\|\|p_{\alpha}\|},\;\frac{\delta}{4N\sqrt{\|x\|C_{y}}\|p_{\alpha}\|}\frac{\delta}{4NC_{y}\|p_{\alpha}\|},\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\frac{\|p_{\alpha}^{3}y_{1}\|}{2\|p_{\alpha}y_{1}\|},\cdots,\frac{\|p_{\alpha}^{3}y_{N}\|}{2\|p_{\alpha}y_{N}\|}\right\}.$
+Choose $F\in\mathcal{F}$ and the families $\{g_{t,1}\}_{t\in F},\dots,\{g_{t,N}\}_{t\in F},\{d_{t}\}_{t\in F}$ satisfying the assumptions in $(2)$ with respect to $\epsilon$ and $p_{\alpha}$.For each $\alpha$ and every $t\in F,l\in\{1,\dots,N\}$ we define the operator $S_{t,l,\alpha}$ on $\mathcal{A}$ by $S_{t,l,\alpha}(a)=\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1})\Phi_{t,l}^{-1}(a)$ for all $a\in\mathcal{A}.$ We observe now that for each $t\in F$ and $l\in\{1,\dots,N\}$ it holds that
+$\parallel T_{t,l}\!\bigl(d_{t}p_{\alpha}x\bigr)\parallel=\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\Phi_{t,l}(p_{\alpha}x)\,\parallel\leq\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\parallel\,\parallel\,\Phi_{t,l}(p_{\alpha}x)\,\parallel$
+$\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l})\,d_{t}\,\parallel\;\parallel\,x\,\parallel\|p_{\alpha}\|$
+and
+$\parallel S_{t,l,\alpha}\!\bigl(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\parallel=\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel$
+$\leq\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}\,g_{t,l})\,\parallel\,\parallel\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel=\parallel\,b_{t,l,\alpha}^{-1}\,g_{t,l}\,\parallel\;\parallel\,y_{\ell}\,\parallel\|p_{\alpha}\|.$
+Moreover, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we have
+$\parallel T_{t,r}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell})\bigr)\parallel=\parallel\,b_{t,r}\,\Phi_{t,r}\,(\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\bigr)\,\Phi_{t,r}\,\bigl(\Phi_{t,l}^{-1}\,(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\bigr)\,\parallel$
+$\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\,(g_{t,l})\,\Phi_{t,l}^{-1}\,(p_{\alpha}\,y_{\ell})\,\parallel$
+$\leq\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}g_{t,l})\,\,\parallel\,\parallel y_{\ell}\parallel\|p_{\alpha}\|.$
+Finally, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it follows that
+$T_{t,l}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell}))=g_{t,l}p_{\alpha}\,y_{\ell}$
+for all $t\in F$ and each $\ell\in\{1,\dots,N\}.$
+Since $\|d_{t}p_{\alpha}x-p_{\alpha}^{3}\,x\|<\|\epsilon\|\|p_{\alpha}x\|<\|p_{\alpha}^{3}\,x\|$, it follows that $d_{t}p_{\alpha}x\neq 0.$ Similarly, $g_{t,l}p_{\alpha}y_{\ell}\neq 0$ for all $t\in F$ and $\ell\in\{1,\dots,N\}$.
+Hence, since by the assumption we have that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0,$ it follows that $T_{t,\ell}(d_{t}p_{\alpha}x)\neq 0$ for each $t\in F$ and $\ell\in\{1,\dots,N\}.$ Similarly, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it folllows that $S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\neq 0$ for all $t\in F$ and $\ell\in{1,\dots,N}.$
+For each $t\in F$, we put
+$x_{t}:=d_{t}p_{\alpha}x+\frac{\sqrt{\sum_{l=1}^{N}\,\parallel T_{t,l}(d_{t}p_{\alpha}x)\parallel}}{\sqrt{\sum_{l=1}^{N}\,\parallel S_{t,l,\alpha}(g_{t,l}p_{\alpha}y_{\ell})\parallel}\qquad}\,{\sum_{l=1}^{N}\,S_{t,r,\alpha}(g_{t,r}p_{\alpha}y_{r})}.$
+By triangle inequality and choice of $\varepsilon$ one can deduce that for each $t\in F$, we have
+$x_{t}\in O,\ \text{and}\ \frac{\sqrt{\sum_{\ell=1}^{N}\|S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\|}}{\sqrt{\sum_{\ell=1}^{N}\|T_{t,\ell}(d_{t}p_{\alpha}x)\|}}T_{t,r}(x_{t})\in V_{r}.$
+for all $r\in\{1,\dots,N\}.$
+We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system is only needed for the proof of the implication $(1)\Rightarrow(2)$.
+Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
+The next corollary can easily be deduced from Theorem 5.1.
+Corollary 5.2.
+Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}.$ If there exist dense subsets $\mathcal{A}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ every fixed $\alpha$ and all $x\in\mathcal{A}_{0},y_{s}\in\mathcal{B}_{s}$ we have
+$\lim_{k\to\infty}\|b_{n_{k},s,\alpha}^{-1}y_{s}\|\|\Phi_{n_{k},l}^{-1}(b_{n_{k},l})x\|=0,$
+and, in addition, for each distinct $s,\ell\in\{1,\ldots,N\}$ it holds that
+$\lim_{k\to\infty}\|\Phi_{n_{k},\ell}^{-1}(b_{n_{k},\ell})\Phi_{n_{k},s}^{-1}(b_{n_{k},s,\alpha}^{-1}y_{s})\|=0,$
+then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\cdots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
+Remark 5.3.
+The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
+Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
+Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
+$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
+be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
+$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case.
+Further, we define the operator
+$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
+for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 5.1 in this setting.
+5.2. Applications
+We will now apply our results to the case of generalized weighted bilateral shift on $\ell_{2}(\mathcal{C}).$ More precisely, put $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
+$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
+The multiplication on $\mathcal{A}_{1}$ is defined similarly component-wise.
+We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
+For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$ and let $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$.
+For each $n\in\mathbb{N}$ and $s\in\{1,\ldots,N\},$ we define the map $\Phi_{n,s}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
+$\Phi_{n,s}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}},$
+for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n,s}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}$ for all $j\in\mathbb{Z}$).
+Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ If $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also, then, by some calculations it can be checked that $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\cdots,\{\Phi_{n,N}\}_{t\in\mathbb{N}}$ is a disjoint aperiodic system with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
+For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\{W_{n,j}^{(\ell)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(\ell)}$ has a bounded inverse.
+Then, for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, we let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
+$(b_{n,\ell})_{i}=W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n},$
+for all $i\in\mathbb{Z}.$ Since for each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\},$ the operator $W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n}$ is invertible, it follows that the condition (C) is satisfied in this case.
+Then we let $T_{n,\ell}$ be the operator on $\mathcal{A}$ given by
+$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a),\qquad\text{for all }a\in\mathcal{A}.$
+Once again, since $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}=\{\tilde{p}_{J,m}^{3}\}_{(J,m)\in\mathbb{N}^{2}}$ is a left approximate unit for $\mathcal{A}$ (so that this left approximate unit is indexed over $\mathbb{N}^{2}$), by some calculations one can check that for all $n,J,m\in\mathbb{N},\,s\in\{1,\dots,N\}$ it holds that
+$(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=\mathcal{U}_{s}^{-r_{s}n}\,W^{(s)\,-1}_{n,i},$
+for all $i\in\mathbb{Z}$ with $J\leq i\leq J,$ and $(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=0$ for all $i\in\mathbb{Z}$ with $i\leq-J-1$ and all $i\in\mathbb{Z}$ with $i\geq J+1.$ In fact, we write just $b_{n,s,J}^{-1}$ instead of $b_{n,s,(J,m)}^{-1}$ because it does only depend on $J$ and not on $m.$
+Finally, note also that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|.$
+Taking all this into account, it is not hard to deduce Theorem 4.2 as a special case of Theorem 5.1.
+From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ In view or Corollary 5.2, by arguments similar to those in the proofs of [27, Theorem 3.2] and [28, Proposition 2.7], it is neither hard to deduce the following proposition in this special case.
+Proposition 5.4.
+Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
+$\lim_{k\to\infty}\big\|W_{n_{k},j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{n_{k},j+r_{\ell}n_{k}}^{(\ell)}\,x^{(j)}\big\|=0,$
+and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
+$\lim_{k\to\infty}\|W_{n_{k},j-r_{s}n_{k}+r_{\ell}n_{k}}^{(\ell)}\,W_{n_{k},j}^{(s)-1}y_{s}^{(j)}\|=0,$
+then the sequences of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
+Next we let $\Omega$ be a locally compact Hausdorff space, $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$, where $C_{b}(\Omega)$ denotes the space of all bounded continuous functions on $\Omega$.
+We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
+$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in C_{b}(\Omega)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|_{\infty}\leq M_{y}\ \forall j\}.$
+The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
+For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
+For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
+[…]
+Then it is easily checked that $\{\tilde{p}_{(K_{1},K_{2},J)}^{3}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}$ satisfies the condition $(\mathcal{P}.)$
+For each $\ell\in\{1,\dots,N\}$ and $k\in\mathbb{N}$, we let $\tilde{\Phi}_{k,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
+$\tilde{\Phi}_{k,\ell}\big(\{f_{j}\}_{j\in\mathbb{Z}}\big)=\{f_{j-k}\circ\alpha_{\ell}^{k}\,\}_{j\in\mathbb{Z}},$
+for all $\{f_{j}\}_{j\in\mathbb{Z}}\subset\mathcal{A}$. Let now $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, set
+$\Phi_{n,\ell}=\tilde{\Phi}_{r_{\ell}n,\ell}.$
+Clearly, the system of isometric isomorphisms $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is disjoint aperiodic with respect to
+$\{\tilde{p}_{(K_{1},K_{2},J)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}.$
+Also, the condition (R) is satisfied in this case.
+For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
+$(b_{n,\ell})_{j+r_{\ell}n}=\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)\quad\text{for all }j\in\mathbb{Z}.$
+Since for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ the function $w_{j}^{(\ell)}>0,$ it follows that the condition (C) is satisfied in this case.
+Now, for each $K_{2}\subseteq\Omega$ compact, we can find some $v_{K_{2}}\in C_{0}(\Omega)$ with $0\leq v_{K_{2}}\leq 1$ such that ${v_{K_{2}}}_{|_{K_{2}}}=1.$ Recalling that the approximate unit for $\mathcal{A}$ in this case is indexed over all triples $(K_{1},K_{2},J)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }J\in\mathbb{N},$ for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, and every fixed triple $(K_{1},K_{2},J)$ we let $b_{n,\ell,(K_{1},K_{2},J)}^{-1}\in\mathcal{A}_{1}$ be given by
+$(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=\frac{v_{K_{2}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)}$
+for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ and $(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=0$ else.
+Then, for all $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ we let $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
+$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a)\quad\text{for all }a\in\mathcal{A}.$
+The next corollary can be deduced by some calculations from Theorem 5.1.
+Corollary 5.5.
+Under the above notation and assumptions, the following statements are equivalent:
+(1) The sequences of operators
+$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
+are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
+(2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
+for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
+In sequel, we assume that $\Omega=\mathbb{R}.$ and $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [26], for $\tau\in C_{b}(\mathbb{R}),$ we put
+$\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
+For each $f\in\mathcal{C}_{\tau}$ we define
+$\|f\|_{\tau}:=\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}.$
+Then, $\mathcal{C}_{\tau}$ is a Banach algebra [23]. We will call this algebra Segal algebra corresponding to $\tau$.
+As observed in [26], for a homeomorphism $\alpha$ of $\mathbb{R},$ we consider corresponding Segal algebra $\mathcal{C}_{\tau\circ\alpha^{k}}$ for all $k\in\mathbb{Z}$.
+Clearly, $f\cdot(\tau\circ\alpha^{k})\in\mathcal{C}$ for all $k\in\mathbb{Z}$ whenever $f\in C_{0}(\mathbb{R})$ since $\tau\circ\alpha^{k}\in C_{b}(\mathbb{R})$ for all $k\in\mathbb{Z}.$ By definition, $\|f\|_{\infty}\leq\|f\|_{\tau\circ\alpha^{k}}$ whenever $f\in\mathcal{C}_{\tau\circ\alpha^{k}}.$ We recall now the following lemmas from [26].
+Lemma 5.6.
+[26, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
+  1. (1)
+    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}=1$ on $K_{\epsilon_{2}}^{(\tau)}$,
+  2. (2)
+    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in C_{c}(\mathbb{R}),$
+  3. (3)
+    ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\subseteq|\tau|^{-1}(\left[0,\epsilon_{1}\right])).$
+Lemma 5.7.
+[26, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
+Lemma 5.8.
+[26, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$.
+Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
+As in [26], let
+${\rm c}_{0}^{\mathcal{C},\tau}:=\Big\{s:=(s_{k})\in\prod_{k\in\mathbb{Z}}\mathcal{C}_{\tau\circ\alpha^{k}}:\,\lim_{k\rightarrow\infty}\|s_{k}\|_{\tau\circ\alpha^{k}}=\lim_{k\rightarrow\infty}\|s_{-k}\|_{\tau\circ\alpha^{-k}}=0\Big\}$
+and we equip ${\rm c_{0}}^{\mathcal{C},\tau}$ with the norm
+$\|s\|_{0}:=\sup_{k\in\mathbb{Z}}\|s_{k}\|_{\tau\circ\alpha^{k}}.$
+Since by [23] we have that $\mathcal{C}_{\tau\circ\alpha^{k}}$ is Banach algebra for all $k\in\mathbb{Z},$ it is not hard to check that ${\rm c}_{0}^{\mathcal{C},\tau}$ is a non-unital Banach algebra.
+We will from now on denote ${\rm c_{0}}^{\mathcal{C},\tau}$ by $\mathcal{A}.$ Moreover, we let $\mathcal{A}_{1}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is straightforward to check that the condition (E) is satisfied also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}$ (and $\mathcal{A}_{1}$ is as above in the previous case).
+For $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1},J\in\mathbb{N},$ and any collection of compact subsets $K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}$ of $\mathbb{R},$ we let $\tilde{p}_{(K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1}}\in\mathcal{A}$ be given by
+[…]
+(where $\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is the function from Lemma 5.6 with compact ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$).
+By the proof of Lemma 5.7 (see [26]) one can deduce that the condition ($\mathcal{P}$) is satisfied in this case.
+Let now $\alpha=\alpha_{1}=\cdots=\alpha_{N}.$ For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\Phi_{n,\ell}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is not hard to check that the system $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is again a system of isometric isomorphisms also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}.$ Moreover, it is disjoint apreiodic and satisfies the condition (R) with respect to
+$\{\tilde{p}_{((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}\}.$
+For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\mathbb{R}),b_{n,\ell}\in\mathcal{A}_{1}$ and $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be as above in the previous case (when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$).
+Now, given $\epsilon_{1}\in(0,1),$ we choose some $\epsilon_{3}\in(\epsilon_{1},1).$ Since ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$ and ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is compact, by Lemma 5.6, we can find some function $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\in\mathcal{C}_{\tau\circ\alpha^{j}}$ such that $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}(t)=1$ for all $t\in supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}},\text{ }supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}$ is compact and $supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\subseteq|\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{3}\right]).$ Therefore, for each $n\in\mathbb{N},\ell\in\{1,\dots,N\},J\in\mathbb{N},\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}$ and any collection of compact subsets
+$K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})},$
+we let $b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1}\in\mathcal{A}_{1}$ be given by
+$(b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1})_{j+r_{\ell}n}=\frac{v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j+r_{\ell}n})},\epsilon_{1}}\bigr),\epsilon_{3}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha^{n-k}\big)}$
+for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ (where we choose $\epsilon_{3}\in(\epsilon_{1},1)$) and $(b_{n,\ell,(\epsilon_{2},\epsilon_{1}J)}^{-1})_{j+r_{\ell}n}=0$ else.
+By some calculations, it is not hard to deduce the next corollary as a corollary of Theorem 5.1.
+Corollary 5.9.
+Under the above notation and assumptions, the following statements are equivalent:
+(1) The sequences of operators
+$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
+are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
+(2) For every $J\in\mathbb{N}$ and finite collection
+$\{K_{\epsilon_{J}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{-J+1}}^{(\tau\circ\alpha^{-J+1})},\ldots,K_{\epsilon_{0}}^{\tau},\ldots,K_{\epsilon_{J}}^{(\tau\circ\alpha^{J})}\}$
+of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
+for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$
+Example 5.10.
+Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
+Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 5.5 and Corollary 5.9 are satisfied in this case.

### v7

@@
-On $\mathcal{F}$-semi-transitive families of operators
+Disjoint supercyclicity in Banach algebras
 Abstract.
-Motivated by the concept of disjoint topological transitivity for supercyclicity given in [32], in this paper, we present the concept of disjoint Furstenberg-semi-transitivity.
-Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space.
-Also, we characterize disjoint Furstenberg-semi-transitive compositions of a right multiplier and a generalized bilateral weighted shift on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space.
-As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers.
-Finally, we illustrate our results with concrete examples.
-We end this paper with an algebraic generalization to the case of non-unital Banach algebras.
+In this paper, we consider operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra and we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras.
+It turns out that generalized weighted bilateral shifts on the standard Hilbert C*-module is just a special case of our theory.
+Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are another special case of our theory.
+Also, we illustrate our results with concrete examples.
 Keywords:
 Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
-1. Introduction
-Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators.
-For example, Hilden and Wallen in [21] proved that any unilateral backward weighted shift is supercyclic.
-Afterwards, Salas in [37] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion.
-Supercyclicity of several kinds of operators has also been studied in for instance [16, 32, 34, 40].
-Moreover, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 9, 17].
-On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 7, 19, 20, 31].
-Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [32] and characterize $\mathcal{F}$-semi-transitivity classes of operators on various Banach spaces.
-Now, the research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [36, 37].
-Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published.
-In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 25, 39].
-Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [24, 29, 30].
-Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [24].
-These adjoints considered in [24] act on the space of Radon measures on a locally compact, non-compact Hausdorff space.
-Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of these published papers.
-All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the adjoints of weighted composition operators.
-In Section 3 in this paper, after presenting the concept of $\mathcal{F}$-semi-transitivity, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm (Theorem 3.1).
-A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2.
-We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [24, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
-After pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [26, 27, 41].
-In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [27], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [27].
-This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [27].
-In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [32], and then, as a continuation and extension of the results in [27], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding $\mathcal{F}$-family is finitely invariant.
-Once again it turns out that a special case of this theory is the case when the $\mathcal{F}$-family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.4 and Proposition 4.6 in this paper.
-Also, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 5.10) where we also illustrate the difference between $\mathcal{F}$-semi-transitive (and thus also supercyclic) generalized weighted bilateral shift operators on Hilbert modules and supercyclic classical weighted bilateral shifts on the sequence space $\ell^{2}.$ In addition, we introduce a new concept of $\mathcal{F}$-C*-semi-transitivity for generalized bilateral shift operators on the standard Hilbert C*-module and we characterize disjoint $\mathcal{F}$-C*-semi-transitive generalized weighted bilateral shifts.
-By our very definition, $\mathcal{F}$-semi-transitivity implies $\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as illustrated in Example 4.9 in this paper.
-Next, hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 35].
-Motivated by those results, in Proposition 4.10 and Corollary 4.11 we provide sufficient conditions for disjoint $\mathcal{F}$-semi-transitivity of compositions of right multipliers and generalized weighted bilateral shift operators (where the corresponding weights are multiplied from the left) on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space and we illustrate these results with concrete example where these sufficient conditions are satisfied (Example 4.12).
-Finally, we conclude this paper with a remark (Remark 4.13) where we explain that our results from Section 4 remain valid also if we instead of a separable Hilbert space consider any standard Hilbert module over a unital C*-algebra.
-Section 5 provides algebraic generalization of Theorem 4.2 in this paper to the case of more general non-unital Banach algebras.
-2. Preliminaries
-Throughout this paper, we let $\Omega$ be a locally compact non-compact Hausdorff space and $\alpha$ be an a homeomorphism of $\Omega.$ As usual, $C_{0}(\Omega)$ denotes the space of all continuous functions on $\Omega$ vanishing at infinity, $C_{b}(\Omega)$ denotes the space of bounded, continuous functions on $\Omega$, whereas $C_{c}(\Omega)$ stands for the set of all continuous, compactly supported functions on $\Omega$.
-Both $C_{0}(\Omega)$ and $C_{b}(\Omega)$ are equipped with the supremum norm.
-Moreover, we let $w$ be a positive bounded function on $\Omega$ such that also $w^{-1}:=\frac{1}{w}\in C_{b}(\Omega)$ and we put then $T_{\alpha,w}$ to be the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega).$ Easily, one can see that by the above assumptions $T_{\alpha,w}$ is well-defined and $\|T_{\alpha,w}\|\leq\|w\|_{\sup}$.
-Since $\frac{1}{w}$ is also bounded, then $T_{\alpha,w}$ is invertible and we have
-$T_{\alpha,w}^{-1}f=\frac{f\circ\alpha^{-1}}{w\circ\alpha^{-1}},\quad(f\in C_{0}(\Omega)).$
-Simply we denote $S_{\alpha,w}:=T_{\alpha,w}^{-1}$.
-By some calculation one can see that for each $n\in\mathbb{N}$ and $f\in C_{0}(\Omega)$,
-$T^{n}_{\alpha,w}f=\left(\prod_{j=0}^{n-1}(w\circ\alpha^{j})\right)\cdot(f\circ\alpha^{n})\text{ (1)}$
-and
-$S^{n}_{\alpha,w}f=\left(\prod_{j=1}^{n}(w\circ\alpha^{-j})\right)^{-1}\cdot(f\circ\alpha^{-n}).\text{ (2)}$
-The adjoint $T_{\alpha,w}^{\ast}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm.
-It is straightforward to check that
-$T_{\alpha,w}^{\ast}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
-for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
-${T_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
-and
-${S_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
-For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$
-In this paper, $\mathcal{F}$ represents a Furstenberg family, which is a specific collection of subsets satisfying the following conditions:
-  1. (1)
-    The empty set $\emptyset$ is not an element of $\mathcal{F}$.
-  2. (2)
-    If $A$ belongs to $\mathcal{F}$ and $B$ contains $A$, then $B$ also belongs to $\mathcal{F}$.
-3. $\mathcal{F}$-semi-transitive cosine operator functions
-Assume that $S$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $S.$
-Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$.
-For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$.
-We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set
-$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C_{z}^{\ast}\ (U_{1})\cap U_{2}\neq\emptyset\}$
-is an element of $\mathcal{F}$.
-The next theorem gives some sufficient conditions for the family $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ to be $\mathcal{F}$-semi-transitive.
-Theorem 3.1.
-We have $(ii)\Rightarrow(i)$.
-  1. $(i)$
-    $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive;
-  2. $(ii)$
-    for each compact subset $K$ of $\Omega$, each $\varepsilon>0$, and any two non-zero measures $\mu$ $\nu\in M(\Omega)$ with $|\mu|(K^{c})=|\nu|(K^{c})=0$, there exists some $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ of $K$ such that
-    $|\mu|(A_{z})<\varepsilon\text{, }|\nu|(A_{z})<\varepsilon\text{, }D_{z}\cap E_{z}=\emptyset\text{, }A^{c}_{z}\cap K=D_{z}\cup E_{z},$
-    for all $z\in F$, and
-    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
-    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
-    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
-    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
-    $\sup_{t\in D_{z}}\left(\prod_{j=0}^{1}\left(w_{z}\circ\alpha_{z}^{j}\right)(t)\right)<\varepsilon,$
-    $\sup_{t\in E_{z}}\left(\prod_{j=1}^{2}\left(w_{z}\circ\alpha_{z}^{-j}\right)^{-1}(t)\right)<\varepsilon\ \ \text{ for all }z\in F.$
-Proof.
-Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$.
-Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$.
-As in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$.
-Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$.
-Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
-$\varepsilon:=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}.$
-For each $z\in F$, let $\tilde{\mu_{z}}$, $\tilde{\nu_{z}}$, $\tilde{\tilde{\nu_{z}}}$ be the elements of $M(\Omega)$ given by $\tilde{\mu_{z}}(B)=\mu(B\cap A_{z}^{c}\cap K)$, $\tilde{\nu_{z}}(B)=\nu(B\cap D_{z})$, and $\tilde{\tilde{\nu_{z}}}(B)=\nu(B\cap E_{z})$ for every $z\in F$ and any Borel subset $B$ of $\Omega$.
-Then
-$|\mu-\tilde{\mu_{z}}|(\Omega)=|\mu|(A_{z})<\varepsilon<\delta\ \ \text{ and }\ \ |\mu-\tilde{\nu_{z}}-\tilde{\tilde{\nu_{z}}}|(\Omega)=|\nu|(A_{z})<\varepsilon<\delta,$
-[…]
-$\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|>0\ \ \text{ and }\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|>0.$
-For each $z\in F$, put
-[…]
-Then
-$\|\phi_{z}-\mu\|\leq\|\mu-\tilde{\mu_{z}}\|+2\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}$
-and
-$\left\|\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})-\nu\right\|$ $\leq\frac{1}{2}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}\ \sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}$
-$+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
-for all $z\in F$. Now, by exactly the same arguments as in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
-$\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
-$\left\|S^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\mu\|\right\}\text{,}$
-$\left\|T^{\ast}_{z}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{w_{z}(t)\|\nu\|\right\}\text{,}$
-$\left\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\nu\|\right\}\text{,}$
-$\left\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{\prod_{j=0}^{1}(w_{z}\circ\alpha_{z}^{j})(t)\|\nu\|\right\}\text{,}$
-and
-$\left\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{\prod_{j=1}^{2}(w_{z}\circ\alpha_{z}^{-j})^{-1}(t)\|\nu\|\right\}\text{.}$
-Utilizing this fact together with assumptions $(ii)$ w.
-r.
-to $\varepsilon=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}$ and the fact that $B(\mu,\delta)\subset U_{1}$ and $B(\nu,\delta)\subset U_{2},$ we deduce that
-$\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})\in U_{1}$
-for all $z\in F$. Hence
-$F\subseteq\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}.$
-Since $\mathcal{F}$ is Furstenberg , we obtain that
-$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}\in\mathcal{F}.$
-Hence, $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive since $U_{1}$ and $U_{2}$ were arbitrary. ∎
-Corollary 3.2.
-Let $w$ be a continuous positive function on $\Omega$ such that $w$ and $\frac{1}{w}$ are bounded and let $\alpha$ be a homeomorphism of $\Omega$.
-Suppose that for each compact subset $K$ of $\Omega$ it holds that
-$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
-and either
-$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{2n-1}\left(w\circ\alpha^{j}\right)(t)\right)=0$
-or
-$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0.$
-Then, for each non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, there exists some $N\in\mathbb{N}$ and some sequence $\{\lambda_{n}\}_{n=N}^{\infty}$ of positive numbers such that $\lambda_{n}C_{n}^{\ast}(U_{1})\cap U_{2}\neq\emptyset\text{ for all }n\geq N,$ where $C_{n}^{\ast}=\frac{1}{2}\left(T^{\ast n}_{\alpha,w}+S^{\ast n}_{\alpha,w}\right)$.
-Proof.
-Let $S=\mathbb{N}$ and $\mathcal{F}$ be the family of all infinite subset of $\mathbb{N}$.
-For each $n\in\mathbb{N}$, we let $\alpha_{n}:=\alpha^{n}$ and $w_{n}:=\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)$.
-Then we consider the case when $A_{n}=E_{n}=\emptyset$ and $D_{n}=K$ (such that $\tilde{\tilde{\nu_{n}}}=0$) for all $n\in\mathbb{N}$, and the case when $A_{n}=D_{n}=\emptyset$ and $E_{n}=K$ for all $n\in\mathbb{N}$, (such that $\tilde{\nu_{n}}=0$ for all $n\in\mathbb{N}$).
-By adapting the proof of Theorem 3.1 for these special cases, we can deduce the statement of corollary.
-∎
-Now we will provide a concrete example where the conditions of Corollary 3.2 are satisfied.
-Example 3.3.
-Let $X=\mathbb{R}$ and $M,\delta$ be positive constants such that $M\geq 2+2\delta$ and $\delta\geq 1.$ Put $\alpha$ to be the function on $\mathbb{R}$ given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}.$ Set
-$w(t)=\begin{cases}M\text{ for }t\leq-1,\\ M+\frac{t+1}{2}(1+\delta-M)\text{ for }t\in[-1,1],\\ 1+\delta\text{ for }t\geq 1.\\ \end{cases}.$
-By some calculations one can check that for each compact subset $K$ of $\Omega$ it holds that
-$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
-and
-$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0,$
-so the conditions of Corollary 3.2 are satisfied.
-However, it is not hard to check by some calculations that for every strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ and every $t\in\mathbb{R}$ it holds that
-$\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
-hence the conditions of [24, Proposition 3.3] are not satisfied.
-This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [24, Proposition 3.3].
-4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
-We start with the following definition.
+1. Introduction and preliminaries
+We recall first the following definition.
 Definition 1.1.
-Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\\ \{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$.
+[18, Definition 4.1] Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$.
 We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
 $\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
-In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$.
-For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
-We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [33, Example 1.3.5].
-Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra.
-Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
-$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
-To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
-$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\|x\|^{2}\sum_{j}y_{j}^{*}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
-For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
-[…]
-As shown in the proof of [27, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
-For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$.
-Further, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ let $\{W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(s)}$ has a bounded inverse.
-Also, assume that $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$.
-Then, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ we let $T_{n,s}$ be the operator on $\ell_{2}(\mathcal{\mathcal{C}})$ defined as
-$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
-By [27], $T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{\mathcal{C}})$ for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\}$.
-In what follows, we will assume that $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
-Now we are ready to give the first result of this section.
-Theorem 4.2.
-Under the above notation and assumptions, the following statements are equivalent.
-(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying
-$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
-and
-$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\},$
-and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
-$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
-Proof.
-$(1)\Rightarrow(2)$ Let $J,m\in\mathbb{N}$ be given.
-Since the families $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-semi-transitive and $\mathcal{F}$ is finitely invariant.
-given $\varepsilon>0$, we can find some $F\subseteq\mathcal{F}$ with $\{1,\dots,2J\}\cap F=\varnothing$ such that for each $n\in F$, there exists some $x_{n}\in\ell_{2}(\mathcal{C})$ and some $\lambda_{n}\in\mathbb{R}^{+}$ with
-$\|(x_{n})-\tilde{p}_{J,m}\|<\varepsilon,$
-$\|\lambda_{n}(T_{n,\ell}x_{n})-\tilde{p}_{J,m}\|<\varepsilon\text{ for all }\ell\in\{1,\dots,N\}.$
-This gives further that
-$\|(x_{n})_{j}-P_{m}\|<\varepsilon,$
-and
-$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon.$
-for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$, because
-$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},$
-and $j+r_{\ell}n>J$ for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $n\in F$.
-Hence
-$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
-$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon$
-for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$. Further, since
-$\|\lambda_{n}(T_{n,s}\,x_{n})_{j}-P_{m}\|<\varepsilon\quad\text{for all }j\in\mathbb{Z}\text{ with }-J\leq j\leq J\text{ and }s\in\{1,\dots,N\},$
-it follows that
-$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}-P_{m}\|<\varepsilon,$
-however,
-$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|=\,\|W_{n,j}^{(s)-1}\frac{1}{\lambda_{n}}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|$
-$\|(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
-once again because $j-r_{s}n<J-2J\leq-J.$ For each $n\in F,$ $\ell,s\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, define
-$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}.$
-Then
-$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
-and
-$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|<\varepsilon,$
-$\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|=\|W_{n,j}^{(s)\,-1}\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|<\varepsilon,$
-so
-$\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
-Finally, if $\ell,s\in\{1,\dots,N\}$ with $\ell\neq s$, then $|r_{\ell}-r_{s}|n>2J$ for all $n\in F$.
-Hence, either $j-r_{s}n+r_{\ell}n<-J$ or $j-r_{s}n+r_{\ell}n>J$ for any $j\in\mathbb{Z}$ with $J\leq j\leq J$.
-Therefore, $(\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}=0,$ hence
-$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=$
-$\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\big\|$
-$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{\ell}^{r_{\ell}n}\big\|$
-$\big\|\lambda_{n}(T_{n,\ell}(x_{n}))_{j-r_{s}n+r_{\ell}n}\big\|=\big\|(\lambda_{n}T_{n,\ell}(x_{n})-\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}\big\|<\varepsilon.$
-for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$, and $s,\ell\in\{1,\dots,N\}$ with $s\neq\ell$.
-$(2)\Rightarrow(1)$ Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$.
-Then there exist some $J,m\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m}x=x$ and $\tilde{p}_{J,m}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}$ since the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
-For each $a\in\ell_{2}(\mathcal{C}),$ let $B_{\delta}(a)$ denote the open ball with centre in $a$ and radius $\delta$ .
-We can find some $\delta>0$ such that $B_{\delta}(x)\subseteq\mathcal{O}$ and $B_{\delta}(y^{(\ell)})\subseteq V_{\ell}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
-$\varepsilon:=\frac{\delta}{2JN(\|x\|+2C_{y}+\delta)}.$
-Choose $F\in\mathcal{F}$ and families
-$\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$
-for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
-For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
-$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
-and
-$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
-Since $\|D_{n,j}-P_{m}\|<\varepsilon$ and $\|G_{n,j}^{(\ell)}-P_{m}\|<\varepsilon<1$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, it follows that $D_{n,j}\neq 0$ and $G_{n,j}^{(\ell)}\neq 0,$ hence $W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\neq 0$, and $W_{n,j}^{(\ell)-1}G_{n,j}^{(\ell)}\neq 0$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$.
-Now we notice that
-$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
-because
-$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
-Moreover, since for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$,
-$(T_{n,\ell}(v_{n}^{(s)}))_{j-r_{s}n+r_{\ell}n}=$
-$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}_{s}^{*\,r_{s}n}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
-it follows that
-$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
-Finally, it is easy to see that by the construction of $v_{n}^{(\ell)}$, we have
-$T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else.}$
-For each $n\in F$, set
-$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
-By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
-$x_{n}\in\mathcal{O}\quad\text{and}\quad\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}T_{n,\ell}(x_{n})\in V_{\ell}$
-for all $n\in F$ and $\ell\in\{1,\dots,N\}$. ∎
-Next, let $\{S_{n}\}_{n\in\mathbb{N}}$ be a sequence of bounded invertible linear operators on $H$ and $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be families of generalized weighted bilateral shifts on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
-For each $n\in F$ and $s\in\{1,\dots,N\}$ we let $S_{n}T_{n,s}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
-$\big(S_{n}T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big):=\{S_{n}W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
-It is not hard to check that the operator $S_{n}T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{C}).$ In fact, $S_{n}T_{n,s}$ is again a generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ with respective weights $\{S_{n}W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ for every $n\in F$ and $s\in\{1,\dots,N\}.$
-We will say that $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-C*-semi-transitive , or shortly $d\mathcal{F}$-C*-semi-transitive , if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $n\in F$ there exists some invertible bounded linear operator $S_{n}$ on $H$ satisfying that
-$\mathcal{O}\cap(S_{n}T_{n,1})^{-1}(V_{1})\cap\dots\cap(S_{n}T_{n,N})^{-1}(V_{N})\neq\varnothing.$
-The next theorem can easily be deduced from Theorem 4.2.
-Theorem 4.3.
-Under the above notation and assumptions, the following statements are equivalent.
-(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in F}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
-$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
-$\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|<\varepsilon,\text{ }\|W_{n,j}^{(s)\,-1}S_{n}^{-1}G_{n,j}^{(s)}\|<\varepsilon,$
-and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
-$\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
-Proof.
-For the proof of $(1)\Rightarrow(2),$ we can proceed in exactly the same way as in the proof of Theorem 4.2 part $(1)\Rightarrow(2)$ by considering $S_{n}$ and $S_{n}^{-1}$ instead of $\lambda_{n}$ and $\tfrac{1}{\lambda_{n}},$ respectively.
-For the proof of $(2)\Rightarrow(1),$ we can also proceed similarly as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$ However, the only difference is that in this case we let now for each $n\in F$, $v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
-$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else,}\end{cases}$
-which clearly gives that
-$\|(v_{n}^{(\ell)})\|\leq\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\|\|y^{(\ell)}\|.$
-Then it suffices to also notice in this case that for each $n\in F$ and $\ell\in\{1,\dots,N\}$ it holds that
-$\|S_{n}T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
-(where $u_{n}$ is the same as in the proof of Theorem 4.2 in the part $(2)\Rightarrow(1)$) and and moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and $n\in F$ we have
-$\|S_{n}T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
-Finally, by the construction of $v_{n}^{(\ell)}$, it follows that
-$S_{n}T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else,}$
-for each $n\in F$ and $\ell\in\{1,\dots,N\}.$
-For each $n\in F$, set
-$x_{n}:=u_{n}+\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
-By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
-$x_{n}\in\mathcal{O}\quad\text{and}\quad S_{n}T_{n,\ell}(x_{n})\in V_{\ell}$
-for all $n\in F$ and $\ell\in\{1,\dots,N\}.$
-Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$ and
-$\{W_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{W_{j}^{(N)}\}_{j\in\mathbb{Z}}\subset B(H)$
-be such that each $W_{j}^{(\ell)}$ $(j\in\mathbb{Z},\,1\leq\ell\leq N)$ is invertible, and such that
-$\|W_{j}^{(\ell)}\|\leq M\quad\text{for all }j\in\mathbb{Z},\,\ell\in\{1,\dots,N\},$
-and some fixed $M>0$. For each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$, we let $W_{n,i}^{(l)}\in B(H)$ be given by
-$W_{n,i}^{(l)}=W_{i}^{(\ell)}W_{i-1}^{(\ell)}\cdots W_{i-r_{\ell}n+1}^{(\ell)}.$
-Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [27]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell}$, i.e.
-[…]
-Then it is not hard to check that for each $\ell\in\{1,\dots,N\}$ it holds that
-[…]
-We will now present an application of Theorem 4.2 and Theorem 4.3 in this concrete case.
-Corollary 4.4.
-Under the above notation and assumptions, the following statements are equivalent.
-(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators in $\mathcal{C}$
-$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
-$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
-$\vdots$
-$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
-such that
-$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
-and
-$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0$
-for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
-$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|=0.$
-Corollary 4.5.
-Under the above notation and assumptions, the following statements are equivalent.
-(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N},$ a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ and sequences of operators in $\mathcal{C}$
-$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
-$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
-$\vdots$
-$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
-such that
-$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
-and
-$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0,$
-$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0,$
-for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$ and satisfying in addition that for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
-$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0.$
-In order to construct concrete examples satisfying the conditions in $(2)$ of Corollary 4.4 and Corollary 4.5, we need first the following technical results.
-Proposition 4.6.
-Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
-$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
-and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
-$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0,$
-then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-Proof.
-The main idea in this proof is motivated by the proofs of [27, Theorem 3.2] and [28, Proposition 2.7].
-Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
-$\{f_{i,j}^{(\eta)}\}_{i\in\mathbb{N}},\subseteq H_{0}^{(j)},\{h_{i,l,j}^{(\eta)}\}_{i\in\mathbb{N}}\subseteq H_{l}^{(j)}\text{ for each }\ell\in\{1,\dots,N\}$
-such that
-$f_{i,j}^{(\eta)}\to e_{\eta},\text{ and }h_{i,\ell,j}^{(\eta)}\to e_{\eta}\quad\text{as }i\to\infty,$
-for all $\eta\in\{-m,\dots,m\}$ and $\ell\in\{1,\dots,N\}$. Since by the assumption we have for all $\ell,s\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ that
-$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}f_{i,j}^{(\eta)}\big\|=0,$
-and for each distinct $s,\ell\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ we have
-$\lim_{k\to\infty}\big\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\dots W_{j-r_{\ell}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\dots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|=0,$
-we can inductively easily construct a subsequence $\{n_{k_{i}}\}_{i\in\mathbb{N}}$ such that
-$\big\|W_{j-r_{s}n_{k_{i}}}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k_{i}}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,f_{i,j}^{(\eta)}\big\|<\tfrac{1}{4m^{2}4^{i}},$
-and
-$\big\|W_{j+r_{\ell}n_{k_{i}}-r_{s}n_{k_{i}}}^{(\ell)}\cdots W_{j-r_{s}n_{k_{i}}+1}^{(\ell)}W_{j-r_{s}n_{k_{i}}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|<\tfrac{1}{2_{m}2^{i}}.$
-For each $i\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$ we define operators $D_{j}^{(i)}$ and $G_{j,\ell}^{(i)}$ on $H$ by
-$(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
-By some calculations and arguments similar to those in the proof of [27, Theorem 3.2] and [28, Proposition 2.7], we can conclude that the sequences of operators
-$\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
-satisfy the conditions in (2) of Corollary 4.4. for every $j\in\mathbb{Z}$ with $J\leq j\leq J$. ∎
-Similarly, based on Theorem 4.3 and Corollary 4.5, we can prove the following proposition.
-Proposition 4.7.
-Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H,$ a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
-$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}\,y_{s}^{(j)}\big\|=0,$
-$\lim_{k\to\infty}\big\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
-and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
-$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}y_{s}^{(j)}\|=0,$
-then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
-Thanks to Proposition 4.6 we will now give an example where we will show that it can happen that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}$ and that at a same time the sequence of operators $\{\tilde{T}^{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$-semi-transitive, which illustrates the important difference between $\mathcal{F}$-semi-transitivity (and thus supercyclicity) of generalized bilateral weighted shifts on Hilbert C*-modules and supercyclicity of the classical bilateral weighted shifts on the sequence space $\ell^{2}.$
-Example 4.8.
-Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$.
-As in [27, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
-Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$.
-For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$.
-If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [27, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
-By our definition, $d\mathcal{F}$-semi-transitivity implies $d\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as the next example illustrates.
-Example 4.9.
-Let again $N=1,r_{1}=1,H=L^{2}(\mathbb{R})$ and $\alpha$ be the function on $\mathbb{R}$ given by $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ Let $v=2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}$ and $\tilde{S}$ be the weighted composition operator on $H$ with respect to $v$ and $\alpha,$ that is $\tilde{S}(f)=v\cdot(f\circ\alpha)$ for all $f\in H.$ For each $n\in\mathbb{N},$ put $S_{n}:=\tilde{S}^{n}.$ If we let for each $j\in\mathbb{Z},$ $W_{j}$ be the identity operator on $H,$ then it is not hard to check that the conditions of Proposition 4.7 are satisfied with $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ Thus, the generalized non-weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ is $\mathcal{F}$-C*-semi-transitive, however, by Theorem 4.2 it can not be $\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
-Hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 35].
-Motivated by those results, in the rest of this paper, we shall study the dynamics of a composition of a right multiplier and generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ defined in the beginning of this section.
-To this end, let $\tilde{W_{1}},\dots,\tilde{W_{N}}$ be invertible, bounded linear operators on $H.$ For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $R_{s,n}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
-$R_{s,n}(\{x_{i}\}_{i\in\mathbb{Z}})=(\{x_{i}\cdot\mathcal{U}_{s}^{*\,r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{i\in\mathbb{Z}})$
-for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [33, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section.
-For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
-$\big((T_{n,s}\circ R_{s,n})(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
-We will now present an extension of Theorem 4.2 in this setting.
-Proposition 4.10.
-Under the above notation and assumptions, we have $(2)\Rightarrow(1).$
-(1) The families of operators $\{T_{n,1}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that we can find families
-$\{\tilde{D_{n}}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},$
-$\{D_{n,-J}\}_{n\in F},\{G_{n,-J}^{(1)}\}_{n\in F},\dots,\{G_{n,-J}^{(N)}\}_{n\in F},$
-$\{D_{n,-J+1}\}_{n\in F},\{G_{n,-J+1}^{(1)}\}_{n\in F},\dots,\{G_{n,-J+1}^{(N)}\}_{n\in F},$
-$\vdots$
-$\{D_{n,J}\}_{n\in F},\{G_{n,J}^{(1)}\}_{n\in F},\dots,\{G_{n,J}^{(N)}\}_{n\in F},$
-of compact operators on $H$ satisfying
-$\|\tilde{D_{n}}-P_{m}\|<\varepsilon,\|D_{n,j}-P_{m}\|<\varepsilon,\|\tilde{G_{n}}^{(s)}-P_{m}\|<\varepsilon,\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
-and
-$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\|<\varepsilon^{2}$
-for all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\};$ and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ it holds that
-$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|<\varepsilon.$
-Proof.
-Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$.
-Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [33, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
-$\varepsilon:=\frac{\delta}{(2JN+3)(\|x\|+2C_{y}+\delta)}.$
-Choose $F\in\mathcal{F}$ and families
-$\{\tilde{D_{n}}\}_{n\in F},\{D_{n,j}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},\{G_{n,j}^{(N)}\}_{n\in F}$
-for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
-For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
-$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j}\tilde{D_{n}},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
-and
-$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
-By similar arguments as in the proof of Theorem 4.2 part $(2)\Rightarrow(1)$, we can deduce that $\tilde{D_{n}}\neq 0$ and $\tilde{G_{n}}^{(s)}\neq 0$ for all $n\in F,s\in\{1,\dots,N\},$ and moreover
-$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|x\|$
-and
-$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|y^{(s)}\|.$
-For each $n\in F$, set
-$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|\|\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
-and proceed further as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$
-Let again, for each $\ell\in\{1,\dots,N\},$ $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell},$ as defined in the beginning of this section.
-Recalling that the involution on $\mathcal{C}$ is an isometry, by similar arguments as in the proof of Proposition 4.6, we easily obtain the following corollary of Proposition 4.10 in the special case when $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$
-Corollary 4.11.
-Under the above notation and assumptions, if there exist dense subsets
-$\tilde{H_{0}},\tilde{H_{1}},\dots,\tilde{H_{N}}$,
-$H_{0}^{(-J)},H_{1}^{(-J)},\dots,H_{N}^{(-J)},$
-$\vdots$
-$H_{0}^{(J)},H_{1}^{(J)},\dots,H_{N}^{(J)}$
-of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ each $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $\tilde{x}\in\tilde{H_{0}},x^{(j)}\in H_{0}^{(j)}$, $\tilde{y_{s}}\in\tilde{H_{s}},y_{s}^{(j)}\in H_{s}^{(j)}$ we have
-$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\|((\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}\tilde{x}\|=0,$
-and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
-$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}(\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|=0,$
-then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
-We end this paper with an example where the conditions of Corollary 4.11 are satisfied.
-Example 4.12.
-Let $N=1$ and $H=L^{2}(\mathbb{R})$.
-For each $j\in\mathbb{Z},$ let $W_{j}$ be the identity operator on $H.$ Put $\tilde{\alpha}$ to be the function on $\mathbb{R}$ given by $\tilde{\alpha}(t)=t-1$ for all $t\in\mathbb{R}$ and $\tilde{w}==2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}.$ If we let $\tilde{W}$ be the weighted composition operator on $H$ with respect to $\tilde{\alpha}$ and $\tilde{w}$, then by some calculations it can be checked that the the conditions of Corollary 4.11 are satisfied with $\tilde{H_{0}}=\tilde{H_{1}}=H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$
-Remark 4.13.
-We notice that due to [33, Proposition 2.2.1] our results in this section remain valid if we instead of a separable Hilbert space $H$ consider any standard Hilbert module $\ell_{2}(\mathcal{A})$ over a unital C*-algebra $\mathcal{A}.$
-5. Algebraic generalizations and axiomatic approach
-We recall first the following definition.
 Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$.
@@
 Also, we will assume that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0.$ Such condition will be called the condition (C) throughout the manuscript.
-5.1. Main results
+2. Main results
 We have the following theorem.
@@
 for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 2.1 in this setting.
-5.2. Applications
-We will now apply our results to the case of generalized weighted bilateral shift on $\ell_{2}(\mathcal{C}).$ More precisely, put $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
+Now we set $N=1$ and we will therefore write $b_{t}$, $\Phi_{t}$ and $T_{t}$ instead of $b_{t,1}$, $\Phi_{t,1}$ and $T_{t,1},$ respectively.
+We will now assume that $b_{t}$ is invertible for each $t\in\mathcal{S}$ and we will consider the map $S_{t}:\mathcal{A}\rightarrow\mathcal{A}$ given by $S_{t}(a)=\Phi_{t}^{-1}(b_{t}^{-1})\,\Phi_{t}^{-1}(a)$ for all $a\in\mathcal{A}$.
+Obviously, $S_{t}=T_{t}^{-1}$ for each $t\in\mathcal{S}.$
+We set $C_{t}:=\frac{1}{2}(T_{t}+S_{t})$ for all $t\in\mathcal{S}$.
+Proposition 2.4.
+Under the above notation and assumptions, we have $(2)\Rightarrow(1)$.
+(1) The family $\{C_{t}\}_{t\in\mathcal{S}}$ is $\mathcal{F}$–semi–transitive.
+(2) For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$, $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ with $e_{t}+f_{t}=g_{t}$ for all $t\in F$ such that
+$\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\qquad\|g_{t}-p_{\alpha}^{2}\|<\varepsilon,$
+$\|\Phi_{t}^{-2}(b_{t})\Phi_{t}^{-1}(b_{t})e_{t}\|<\varepsilon,$
+$\|\Phi_{t}(b_{t}^{-1})\,b_{t}^{-1}\,f_{t}\|<\varepsilon,$
+$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
+$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|b_{t}^{-1}f_{t}\|<\varepsilon^{2},$
+$\|b_{t}^{-1}\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
+$\|b_{t}^{-1}\,d_{t}\|\,\|b_{t}^{-1}\,f_{t}\|<\varepsilon^{2},\text{ for all }t\in F.$
+Proof.
+Given non–empty open subsets $\mathcal{O}$ and $\mathcal{V}$ of $\mathcal{A}$, as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can find some $\delta>0$ and some $p_{\alpha}^{3}x\in\mathcal{O}\setminus\{0\}$, $p_{\alpha}^{3}y\in\mathcal{V}\setminus\{0\}$ such that $\delta$–neighbourhoods of $p_{\alpha}^{3}x$ and $p_{\alpha}^{3}y$ are contained in $\mathcal{O}\setminus\{0\}$ and $V\setminus\{0\}$, respectively.
+Set
+$\varepsilon:=\min\left\{\frac{\delta}{4\|p_{\alpha}\|(\|x\|+\|y\|)},\;\frac{\delta}{16\|p_{\alpha}\|\sqrt{(\|x\|\|y\|)}},\;\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\;\frac{\|p_{\alpha}^{3}y\|}{2\|p_{\alpha}y\|}\right\}.$
+Choose $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$ , $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ satisfying the assumptions in (2) with respect to $p_{\alpha}$ and $\varepsilon$.
+Since $e_{t}+f_{t}=g_{t}$ and $g_{t}\,p_{\alpha}y\neq 0$ (as shown in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$), it follows that either $e_{t}\,p_{\alpha}y\neq 0$ or $f_{t}\,p_{\alpha}y\neq 0$.
+Then, by the same arguments as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can show that either $T_{t}(e_{t},p_{\alpha}y)\neq 0$ or $S_{t}(f_{t},p_{\alpha}y)\neq 0$.
+Further, we notice that for each $t\in F$ it holds that
+$\|T_{t}^{2}(e_{t},p_{\alpha}y)\|=\|b_{t}\,\Phi_{t}\,(b_{t})\,\Phi_{t}^{2}\,(e_{t}\,p_{\alpha}y)\|$
+$\leq\|\Phi_{t}^{-2}\,(b_{t})\,\Phi_{t}^{-1}\,(b_{t})\,e_{t}\|\,\|p_{\alpha}\|\,\|y\|$
+and,
+$\|S_{t}^{2}(f_{t},p_{\alpha}y)\|=\|\Phi_{t}^{-1}(b_{t}^{-1})\cdot\Phi_{t}^{-2}(b_{t}^{-1}f_{t}\,p_{\alpha}y)\|\leq\|\Phi_{t}(b_{t}^{-1})b_{t}^{-1}f_{t}\|\,\|p_{\alpha}\|\,\|\|y\|.$
+For each $t\in F,$ put
+$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}+p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
+By the above inequalities and inequalities in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$, the choice of $\varepsilon$ and the triangle inequality, we can deduce that $v_{t}\in\mathcal{O}$ and
+$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}+p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
+By putting either $e_{t}=0$ (so that $f_{t}=g_{t}$), or by putting $f_{t}=0$ (so that $e_{t}=g_{t}$) in Proposition 2.4, then we can get the following corollary.
+Corollary 2.5.
+Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$.
+If there exist dense subsets $\mathcal{B}_{1},\mathcal{B}_{2}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ such that for every fixed $p_{\alpha}$ it holds that either and
+$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-2}(b_{n_{k}})\,\Phi_{n_{k}}^{-1}(b_{n_{k}})\,y\|=\lim_{k\to\infty}\|b^{-1}_{n_{k}}x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=$
+$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=0$
+or
+$\lim_{k\to\infty}\|\Phi_{n_{k}}(b_{n_{k}}^{-1})b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|b_{n_{k}}^{-1}x)\|\cdot\|b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})\|\|b_{n_{k}}^{-1}y\|=0$
+for all $x\in\mathcal{B}_{1}$ and $y\in\mathcal{B}_{2}$, then the family $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive.
+3. Applications
+From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$.
+For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
+Moreover, we will denote by $\Omega$ be a locally compact Hausdorff, non-compact space.
+As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity.
+Both these spaces will be equipped with supremum norm.
+In addition, we let $C_{c}(\Omega)$ is the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
+3.1. Applications to generalized weighted shifts on Hilbert C*-modules
+In this subsection, we will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [21, Example 1.3.5].
+Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra.
+Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
+$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
+To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
+$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
+For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
+[…]
+As shown in the proof of [16, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
+We put then $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
 $\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
@@
 We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
-For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$ and let $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$.
-For each $n\in\mathbb{N}$ and $s\in\{1,\ldots,N\},$ we define the map $\Phi_{n,s}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
-$\Phi_{n,s}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}},$
-for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n,s}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}$ for all $j\in\mathbb{Z}$).
-Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ If $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also, then, by some calculations it can be checked that $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\cdots,\{\Phi_{n,N}\}_{t\in\mathbb{N}}$ is a disjoint aperiodic system with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
-For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\{W_{n,j}^{(\ell)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(\ell)}$ has a bounded inverse.
-Then, for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, we let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
-$(b_{n,\ell})_{i}=W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n},$
-for all $i\in\mathbb{Z}.$ Since for each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\},$ the operator $W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n}$ is invertible, it follows that the condition (C) is satisfied in this case.
-Then we let $T_{n,\ell}$ be the operator on $\mathcal{A}$ given by
-$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a),\qquad\text{for all }a\in\mathcal{A}.$
-Once again, since $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}=\{\tilde{p}_{J,m}^{3}\}_{(J,m)\in\mathbb{N}^{2}}$ is a left approximate unit for $\mathcal{A}$ (so that this left approximate unit is indexed over $\mathbb{N}^{2}$), by some calculations one can check that for all $n,J,m\in\mathbb{N},\,s\in\{1,\dots,N\}$ it holds that
-$(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=\mathcal{U}_{s}^{-r_{s}n}\,W^{(s)\,-1}_{n,i},$
-for all $i\in\mathbb{Z}$ with $J\leq i\leq J,$ and $(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=0$ for all $i\in\mathbb{Z}$ with $i\leq-J-1$ and all $i\in\mathbb{Z}$ with $i\geq J+1.$ In fact, we write just $b_{n,s,J}^{-1}$ instead of $b_{n,s,(J,m)}^{-1}$ because it does only depend on $J$ and not on $m.$
-Finally, note also that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|.$
-Taking all this into account, it is not hard to deduce Theorem 4.2 as a special case of Theorem 5.1.
-From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ In view or Corollary 5.2, by arguments similar to those in the proofs of [27, Theorem 3.2] and [28, Proposition 2.7], it is neither hard to deduce the following proposition in this special case.
+Let now $\mathcal{U}$ be a unitary operator on $H$. For each $n\in\mathbb{N}$ we define the map $\Phi_{n}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
+$\Phi_{n}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}\}_{j\in\mathbb{Z}},$
+for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}$ for all $j\in\mathbb{Z}$).
+Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ Moreover, by some calculations it can be checked that $\{\Phi_{n}\}_{n\in\mathbb{N}}$ is aperiodic with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
+For each $n\in\mathbb{N},$ let $\{W_{n,j}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm such that each $W_{n,j}$ has a bounded inverse and such that $\{W_{n,j}^{-1}\}_{j\in\mathbb{Z}}$ is also uniformly bounded in the norm.
+Then, for each $n\in\mathbb{N}$ we let $b_{n}\in\mathcal{A}_{1}$ be given by
+$(b_{n})_{i}=W_{n,i}\mathcal{U}^{n},$
+for all $i\in\mathbb{Z}.$ Obviously, $b_{n}$ is invertible in $\mathcal{A}_{1}$ for each $n\in\mathbb{N}$ and
+$(b_{n}^{-1})_{i}=\mathcal{U}^{-n}\,W^{\,-1}_{n,i},$
+for all $i\in\mathbb{Z},$ hence, the condition $(C)$ is satisfied in this case.
+Finally, for each $n\in\mathbb{N}$ we let $T_{n}$ be the operator on $\mathcal{A}$ given by
+$T_{n}(a)=b_{n}\,\Phi_{n}(a),\qquad\text{for all }a\in\mathcal{A}.$
+Noticing that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|$, in view of Corollary 2.5 and by arguments similar to those in the proofs of [16, Theorem 3.2] and [17, Proposition 2.7] it is not hard to deduce the following proposition in this special case.
 Proposition 3.1.
-Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
-$\lim_{k\to\infty}\big\|W_{n_{k},j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{n_{k},j+r_{\ell}n_{k}}^{(\ell)}\,x^{(j)}\big\|=0,$
-and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
-$\lim_{k\to\infty}\|W_{n_{k},j-r_{s}n_{k}+r_{\ell}n_{k}}^{(\ell)}\,W_{n_{k},j}^{(s)-1}y_{s}^{(j)}\|=0,$
-then the sequences of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
-Next we let $\Omega$ be a locally compact Hausdorff space, $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$, where $C_{b}(\Omega)$ denotes the space of all bounded continuous functions on $\Omega$.
+Under the above notation and assumptions, if there exist dense subsets $H_{0}$ and $H_{1}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}$ of natural numbers such that either
+$\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W_{n_{k,j+n_{k}}}y\|=\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W_{n_{k},j+n_{k}}y\|$
+$\lim_{k\to\infty}\|W_{n_{k},j+2n_{k}}W_{n_{k},j+n_{k}}y\|=0,$
+or
+$\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W^{-1}_{n_{k},j}y\|=\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W^{-1}_{n_{k},j}y\|=$
+$\lim_{k\to\infty}\|W^{-1}_{n_{k},j-n_{k}}W^{-1}_{n_{k},j}y\|=0,$
+for all $x\in H_{0}$ and $y\in H_{1}$, then $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive on $\mathcal{A}$.
+Example 3.2.
+Let $H=L^{2}(\mathbb{R})$.
+As in [16, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel,\parallel w_{j}^{-1}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
+Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq K$ for all $j\in\mathbb{Z}$ and some $K>0$.
+For each $j\in\mathbb{Z},$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R},$ and let then $V_{j}\in B(H)$ be given by
+$V_{j}(f)=w_{j}\cdot(f\circ\alpha_{j})$
+for all $f\in H.$ For each $n\in\mathbb{N},i\in\mathbb{Z}$, let $W_{n,i}\in B(H)$ be given by
+$W_{n,i}:=V_{i}V_{i-1}\cdots V_{i-n+1}.$
+If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [16, Example 3.3] one can show that the conditions of Proposition 3.1 are satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R})$ in this case.
+On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 3.1 will be satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R}).$
+Next, for $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}.$
+Let $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$.
 We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
@@
 The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
-For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
 For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
@@
 (2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
 for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
 $\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
-In sequel, we assume that $\Omega=\mathbb{R}.$ and $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [26], for $\tau\in C_{b}(\mathbb{R}),$ we put
+Example 3.4.
+Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
+Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 3.3 are satisfied in this case.
+3.2. Applications to generalized weighted composition operators on spaces of operator-valued continuous functions
+In this subsection, we let $C_{0}(\Omega,\mathcal{C})$ denote the non-unital normed algebra of all continuous $\mathcal{C}$-valued functions on $\Omega$ vanishing at infinity and equipped with the norm
+$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|$
+and $C_{b}(\Omega,B(H))$ be the unital normed algebra consisting of all uniformly bounded $B(H)$-valued continuous functions on $\Omega$ equipped with the same norm $\|\cdot\|_{\infty}$.
+Obviously, $C_{0}(\Omega,\mathcal{C})$ is (left) ideal in $C_{b}(\Omega,B(H))$ and the condition $(E)$ is satisfied.
+We will therefore denote $C_{0}(\Omega,\mathcal{C})$ and $C_{b}(\Omega,B(H))$ by $\mathcal{A}$ and $A_{1},$ respectively.
+For every triple $(K_{1},K_{2},m)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }m\in\mathbb{N},$ let $\tilde{P}_{(K_{1},K_{2},m)}\in C_{0}(\Omega,\mathcal{C})$ be given by
+$\tilde{P}_{(K_{1},K_{2},m)}(x)=u_{(K_{1},K_{2})}(x)P_{m}\quad\text{for all }x\in\Omega.$
+We have the following lemma.
+Lemma 3.5.
+The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\in\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
+Proof.
+Let $O$ be a non-empty open subset of $\mathcal{A}$ and $f\in O$.
+Then there exists some $\varepsilon>0$ such that $\varepsilon$– neighbourhood of $f$ is contained in $O$ as well.
+Since $f\in\mathcal{A},$ there exists some compact subset $K_{1}$ of $\Omega$ such that $\|f(x)\|<\frac{\varepsilon}{2}$ for all $x\in\Omega\setminus K_{1}$.
+Now, by continuity of $f,$ for each $x\in K_{1}$ we can find some open set $O_{x}$ of $\Omega$ containing $x$ such that $\|f(u)-f(x)\|<\frac{\varepsilon}{2}$ whenever $y\in O_{x}$.
+Then for all $m\in\mathbb{N}$ and each $y\in O_{x}$ we obtain
+$\|(I-P_{m})f(y)\|\leq\|(I-P_{m})(f(y)-f(x))\|+\|(I-P_{m})f(x)\|$
+$\leq\|f(y)-f(x)\|+\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}+\|(I-P_{m})f(x)\|$
+By [21, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
+The collection $\{O_{x}\}_{x\in K_{1}}$ is an open cover of $K_{1}$, hence there exist a finite subcover $\{O_{x_{1}},\ldots,O_{x_{M}}\}.$ Let $\tilde{m}:=\max\{m_{x_{1}},\ldots,m_{x_{M}}\}.$ Then, by the above we get
+$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)-f(x)\|=\|(I-P_{m})f(x)\|<\varepsilon$ for all $x\in K_{1}$ and $m\geq\tilde{m}$. If $x\in\Omega\setminus K_{1},$ then for all $m\in\mathbb{N}$ we get
+$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)\|+\|f(x)\|=\|u^{3}_{(K_{1},K_{2})}(x)P_{m}f(x)\|+\|f(x)\|\leq 2\|f(x)\|<2\frac{\varepsilon}{2}=\varepsilon.$
+Thus,
+$\|\tilde{P}^{3}_{(K_{1},K_{2},\tilde{m})}f-f\|_{\infty}<\varepsilon,\text{ so }\tilde{P}^{3}_{(K_{1},K_{2},m)}f\in O.$
+Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ We will assume that for every compact subset $K$ of $X$ and every $S\in\mathcal{F}$ there exists some $F\in\mathcal{F}$ with $F\subseteq S$ such that
+$K\cap\alpha_{n,\ell}(K)=\varnothing\quad\text{and}\quad\alpha_{n,\ell}^{-1}(\alpha_{n,s}(K))\cap K=\varnothing$
+for all $n\in F$ and each distinct $s,\ell\in\{1,\dots,N\}.$ Also, we consider families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
+$\Phi_{n,l}:\mathcal{A}_{1}\rightarrow\mathcal{A}_{1}\text{ be given by }\Phi_{t,l}(f):=U_{n,l}^{*}(f\circ\alpha_{n,l})U_{n,l}$
+for all $f\in\mathcal{A}.$ Obviously the system $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\dots,\{\Phi_{n,N}\}_{n\in\mathbb{N}}$ is a system of isometric isomorphisms on $\mathcal{A}$ which is disjoint aperiodic with respect to $\{\tilde{P}_{(K_{1},K_{2},m)}\}$ and, clearly, the condition ($R$) is satisfied.
+Let $\{w_{n,1}\}_{n\in\mathbb{N}},\dots,\{w_{n,N}\}_{n\in\mathbb{N}}$ be families in $\mathcal{A}_{1}$ such that $w_{n,l}(x)$ is invertible for all $x\in\Omega$, $n\in\mathbb{N}$, $l\in\{1,\ldots,N\}$, and such that $\{w^{-1}_{n,1}\}_{n\in\mathbb{N}},\dots,\{w^{-1}_{n,N}\}_{n\in\mathbb{N}}$ also belong to $\mathcal{A}_{1}$ i.e.
+$\sup_{x\in\Omega}\|w^{-1}_{n,l}(x)\|<\infty\quad\text{for all }n\in\mathbb{N}\text{ and }l\in\{1,\ldots,N\}.$
+For each $n\in\mathbb{N}$, and $l\in\{1,\ldots,N\},$ let then $b_{n,l}:=w_{n,l}U_{n,l}$.
+Since each $b_{n,l}$ is invertible in $\mathcal{A}_{1}$, it follows that the condition $(C)$ is satisfied in this case.
+Corollary 3.6.
+Under the above notation and assumptions, if there exist dense subsets $H_{0},H_{1},\ldots,H_{N}$ of $H$ such that for each $s,l\in\{1,\ldots,N\}$ and every compact subset $K$ of $\Omega$ we have that
+$\lim_{n\to\infty}\left[\left(\sup_{x\in K}\|w_{n,s}^{-1}(x)h_{s}\|\right)\cdot\left(\sup_{y\in K}\|(w_{n,\ell}\circ\alpha_{n,\ell}^{-1})(y)h_{0}\|\right)\right]=0$
+for every $h_{s}\in H_{s}$ and $h_{0}\in H_{0},$ and, in addition, for each distinct $s,l\in\{1,\ldots,N\}$ and every $h_{s}\in H_{s}$ we have
+$\lim_{n\to\infty}\left(\sup_{x\in K}\|(w_{n,l}\circ\alpha_{n,l}^{-1}\circ\alpha_{n,s})(x)\,w_{n,s}^{-1}(x)\,h_{s}\|\right)=0,$
+then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\ldots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
+Proof.
+For each $s\in\{1,\ldots,N\}$, the set consisting of elements of the form
+$u_{(K_{1},K_{2})}G_{s,m},$
+where $G_{s,m}\in\mathcal{C}$ is such that
+$G_{s,m}e_{i}\in H_{s}\text{ for all }j\in\{-m,\ldots,m\}\text{ and }G_{s,m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\ldots,m\}$
+is obviously dense in $\mathcal{A}$ by Lemma 3.5 and the density of $H_{s}$ in $H$.
+(Here of course $m$ runs through the whole $\mathbb{N}$ and $(K_{1},K_{2})$ runs through all pairs of compact subsets $K_{1},K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}.$) Let us denote this set by $\mathcal{A}_{s}$ for each $s\in\{1,\ldots,N\}$.
+Similarly, we let $\mathcal{A}_{0}$ be the set consisting of all elements of the form $u_{(K_{1},K_{2})}D_{m}$ where $D_{m}\in\mathcal{C}$ is such that
+$D_{m}e_{j}\in H_{0}\text{ for all }j\in\{-m,\dots m\}\quad\text{and}\quad D_{m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\dots m\}.$
+By some calculations similar to those in the proof of [16, Theorem 3.2] and [17, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 2.2 are satisfied. ∎
+Example 3.7.
+[…]
+3.3. Applications to generalized weighted shifts on function sequence spaces
+In this subsection, we assume that $\Omega=\mathbb{R}.$ and we let now $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [15], for $\tau\in C_{b}(\mathbb{R}),$ we put
 $\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
@@
 Lemma 3.8.
-[26, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
+[15, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
   1. (1)
@@
 Lemma 3.9.
-[26, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
+[15, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
 Lemma 3.10.
-[26, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$.
+[15, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$.
 Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
@@
 of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
+$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
 for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
 $\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$
-Example 5.10.
-Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
-Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 5.5 and Corollary 5.9 are satisfied in this case.

### v8

@@
-Disjoint supercyclicity in Banach algebras
+Disjoint $\mathcal{F}$-semi-transitivity in Banach modules
 Abstract.
-In this paper, we consider operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra and we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras.
-It turns out that generalized weighted bilateral shifts on the standard Hilbert C*-module is just a special case of our theory.
-Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are another special case of our theory.
-Also, we illustrate our results with concrete examples.
+Motivated by the concept of disjoint topological transitivity for supercyclicity given in [28], in this paper, we consider the concept of disjoint Furstenberg-semi-transitivity for operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra.
+Thus, we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras.
+Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are a special case of our theory.
+At the end of the paper, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators acting on the weighted space of Radon measures.
 Keywords:
 Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
-1. Introduction and preliminaries
-We recall first the following definition.
+1. Introduction
+Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators.
+For example, Hilden and Wallen in [17] proved that any unilateral backward weighted shift is supercyclic.
+Afterwards, Salas in [33] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion.
+Supercyclicity of several kinds of operators has also been studied in for instance [12, 28, 30, 36].
+Moreover, quite recently, a characterization of supercyclic and topologically semi-transitive weighted composition operators on solid Banach function spaces, in particular Morrey spaces, and Segal algebras have been given in [24].
+Further, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 6, 13].
+On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 5, 15, 16, 27].
+Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity (Definition 2.1), as a generalization of the concept of topological transitivity for supercyclicity given in [28] and we characterize disjoint $\mathcal{F}$-semi-transitive operators on a large class of non-unital normed algebra (Theorem 3.1 and Corollary 3.2).
+The operators we study are in fact a composition of an isometric isomorphism and a left multiplier on a non-unital normed algebra.
+The reason for considering these operators is because many known classes of operators that have earlier been studied in connection with the dynamics, such as bilateral weighted shifts on $l^{2}(\mathbb{Z}),$ weighted composition operators on algebras of functions, generalized weighted shifts on Hilbert C*-modules and so on, turn out to be of this kind.
+Next, the dynamics of cosine operator functions has been studied in several papers, see [7, 8, 9, 10, 11, 35, 20, 23, 25, 26, 38].
+Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of the published papers, as far as we know.
+All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the above mentioned operators on normed non-unital algebras, i.e.
+the operators that are a composition of an isometric isomorphism and a left multiplier.
+In Proposition 3.4 and Corollary 3.5 we provide some sufficient conditions for the cosine operators functions generated by these operators to be $\mathcal{F}$-semi-transitive on the respective normed algebra.
+Now, an attempt to axiomatically study hypercyclicity of operators on C*-algebras has been done in [20, Section 4].
+This approach is limited to C*-algebras and does not cover many important and relevant cases in linear dynamics.
+The approach in Section 3 in this paper is more general because we consider a large class of non-unital normed algebras that are not necessarily C*-algebras.
+Our approach covers also the case of generalized bilateral weighted shift operators on standard Hilbert C*-modules defined in [19, 20].
+In Section 4 we provide a new application in the context of generalized weighted composition operators on the space of operator valued continuous functions.
+More precisely, we denote by $\mathcal{C}$ the C*-algebra of compact operators on a separable Hilbert space $H,$ and in Corollary 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted composition operators on the algebra $C_{0}(\Omega,\mathcal{C})$ equipped with the supremum-norm, i.e.
+the normed algebra of all $\mathcal{C}$-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space $\Omega.$ The generalized weighted composition operators that we study in this context are in fact a composition of a weighted composition operator on $C_{0}(\Omega,\mathcal{C})$ and a right multiplication by a unitary operator on $H.$ Thus, these operators are at a same time a generalization of (ordinary) weighted composition operators on the classical space $C_{0}(\Omega),$ and elementary (wedge) operators on $\mathcal{C}$ considered in [21].
+We give also a concrete example where the conditions of Corollary 4.2 are satisfied (Example 4.3).
+In Section 5, we study the dynamics of adjoints of weighted composition operators.
+These adjoints act on the weighted space of Radon measures on a locally compact non-compact Hausdorff space.
+The motivation for considering the dynamics of operators on weighted spaces of Radon measures comes from the recent paper [22] where the dynamics of weighted composition operators on weighted solid Banach function spaces have been studied.
+In Theorem 5.5 we characterize the adjoints of weighted composition operators that have a dense set of periodic elements, obtaining in this way an improvement of [23, Proposition 10.3] by deducing equivalent conditions that are weaker than sufficient conditions given in [23, Proposition 10.3] for chaos of the adjoints of weighted composition operators.
+Next, in Theorem 5.6, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators on weighted spaces of Radon measures.
+2. Preliminaries
+We need first the following definition.
 Definition 2.1.
-[18, Definition 4.1] Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$.
+Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and
+$\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$
+be families of bounded linear operators on a Banach space $X$.
 We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
 $\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
-Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$.
+Definition 2.2.
+Let $\mathcal{X}$ be a Banach space, and $(T_{n})_{n\in\mathbb{N}_{0}}$ be a sequence of operators in $B(\mathcal{X})$.
+A vector $x\in\mathcal{X}$ is called a periodic element of $(T_{n})_{n\in\mathbb{N}_{0}}$ if there exists a constant $N\in\mathbb{N}$ such that for each $k\in\mathbb{N}$, $T_{kN}x=x$.
+The set of all periodic elements of $(T_{n})_{n\in\mathbb{N}_{0}}$ is denoted by ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$.
+The sequence $(T_{n})_{n\in\mathbb{N}_{0}}$ is called chaotic if $(T_{n})_{n\in\mathbb{N}_{0}}$ is topologically transitive and ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$ is dense in $\mathcal{X}$}.
+An operator $T\in B(\mathcal{X})$ is called chaotic if the sequence $\{T^{n}\}_{n\in\mathbb{N}_{0}}$ is chaotic.
+Let now $\mathcal{A}$ be a non-unital normed algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$.
 We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$.
@@
 for all $t\in F$ and $r\in\{1,\dots,N\}$.
-For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$
-Then $\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
-$\text{Similarly, for each }t\in F\text{ and }l\in\{1,\dots,N\},\text{ we have}$
+For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$ Then
+$\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon$
+and
+$\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
+Similarly, for each $t\in F$ and $l\in\{1,\dots,N\},$ we have
 $\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})\,\Phi_{t,l}(p_{\alpha})\|\;=\;\|(\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})-p_{\alpha})\,\Phi_{t,l}(p_{\alpha})\|\;\leq\;\|\lambda_{t}\,T_{t,l}(a_{t})-p_{\alpha}\|\;<\;\varepsilon,$
-[…]
-$\text{ set }d_{t}=a_{t}p_{\alpha}.$
-$\text{Then }\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
+and, as shown above,
+$\|a_{t}p_{\alpha}-p_{\alpha}^{2}\|\;<\;\varepsilon$
+For each $t\in F$ and $l\in\{1,\dots,N\},$ set $d_{t}=a_{t}p_{\alpha}.$ Then
+$\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon$
+and
+$|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
 Hence, for all $t\in F$ and $r,l\in\{1,\dots,N\},$ we have
@@
 for all $r\in\{1,\dots,N\}.$
-We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system is only needed for the proof of the implication $(1)\Rightarrow(2)$.
-Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
 The next corollary can easily be deduced from Theorem 3.1.
@@
 Remark 3.3.
-The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
-Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
-Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
-$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
-be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
-$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case.
-Further, we define the operator
-$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
-for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 2.1 in this setting.
+We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system in Theorem 3.1 is only needed for the proof of the implication $(1)\Rightarrow(2)$.
+Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
+Hence, the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is disjoint aperiodic and the condition (R) are not needed in Corollary 3.2.
 Now we set $N=1$ and we will therefore write $b_{t}$, $\Phi_{t}$ and $T_{t}$ instead of $b_{t,1}$, $\Phi_{t,1}$ and $T_{t,1},$ respectively.
@@
 For each $t\in F,$ put
-$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}+p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
+$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
 By the above inequalities and inequalities in the proof of Theorem 3.1 part $(2)\Rightarrow(1)$, the choice of $\varepsilon$ and the triangle inequality, we can deduce that $v_{t}\in\mathcal{O}$ and
-$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}+p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
+$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
 By putting either $e_{t}=0$ (so that $f_{t}=g_{t}$), or by putting $f_{t}=0$ (so that $e_{t}=g_{t}$) in Proposition 3.4, then we can get the following corollary.
@@
 for all $x\in\mathcal{B}_{1}$ and $y\in\mathcal{B}_{2}$, then the family $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive.
+Remark 3.6.
+The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
+Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
+Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
+$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
+be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
+$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case.
+Further, we define the operator
+$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
+for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 3.1 in this setting.
 4. Applications
-From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$.
+From now on and in the rest of this section we will assume that $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$.
 For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
-Moreover, we will denote by $\Omega$ be a locally compact Hausdorff, non-compact space.
+Moreover, we will denote by $\Omega$ be a locally compact, non-compact Hausdorff, space.
 As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity.
 Both these spaces will be equipped with supremum norm.
-In addition, we let $C_{c}(\Omega)$ is the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
-3.1. Applications to generalized weighted shifts on Hilbert C*-modules
-In this subsection, we will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [21, Example 1.3.5].
-Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra.
-Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
-$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
-To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
-$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
-For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
-[…]
-As shown in the proof of [16, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
-We put then $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
-$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
-The multiplication on $\mathcal{A}_{1}$ is defined similarly component-wise.
-We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
-Let now $\mathcal{U}$ be a unitary operator on $H$. For each $n\in\mathbb{N}$ we define the map $\Phi_{n}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
-$\Phi_{n}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}\}_{j\in\mathbb{Z}},$
-for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}$ for all $j\in\mathbb{Z}$).
-Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ Moreover, by some calculations it can be checked that $\{\Phi_{n}\}_{n\in\mathbb{N}}$ is aperiodic with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
-For each $n\in\mathbb{N},$ let $\{W_{n,j}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm such that each $W_{n,j}$ has a bounded inverse and such that $\{W_{n,j}^{-1}\}_{j\in\mathbb{Z}}$ is also uniformly bounded in the norm.
-Then, for each $n\in\mathbb{N}$ we let $b_{n}\in\mathcal{A}_{1}$ be given by
-$(b_{n})_{i}=W_{n,i}\mathcal{U}^{n},$
-for all $i\in\mathbb{Z}.$ Obviously, $b_{n}$ is invertible in $\mathcal{A}_{1}$ for each $n\in\mathbb{N}$ and
-$(b_{n}^{-1})_{i}=\mathcal{U}^{-n}\,W^{\,-1}_{n,i},$
-for all $i\in\mathbb{Z},$ hence, the condition $(C)$ is satisfied in this case.
-Finally, for each $n\in\mathbb{N}$ we let $T_{n}$ be the operator on $\mathcal{A}$ given by
-$T_{n}(a)=b_{n}\,\Phi_{n}(a),\qquad\text{for all }a\in\mathcal{A}.$
-Noticing that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|$, in view of Corollary 2.5 and by arguments similar to those in the proofs of [16, Theorem 3.2] and [17, Proposition 2.7] it is not hard to deduce the following proposition in this special case.
-Proposition 3.1.
-Under the above notation and assumptions, if there exist dense subsets $H_{0}$ and $H_{1}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}$ of natural numbers such that either
-$\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W_{n_{k,j+n_{k}}}y\|=\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W_{n_{k},j+n_{k}}y\|$
-$\lim_{k\to\infty}\|W_{n_{k},j+2n_{k}}W_{n_{k},j+n_{k}}y\|=0,$
-or
-$\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W^{-1}_{n_{k},j}y\|=\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W^{-1}_{n_{k},j}y\|=$
-$\lim_{k\to\infty}\|W^{-1}_{n_{k},j-n_{k}}W^{-1}_{n_{k},j}y\|=0,$
-for all $x\in H_{0}$ and $y\in H_{1}$, then $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive on $\mathcal{A}$.
-Example 3.2.
-Let $H=L^{2}(\mathbb{R})$.
-As in [16, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel,\parallel w_{j}^{-1}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
-Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq K$ for all $j\in\mathbb{Z}$ and some $K>0$.
-For each $j\in\mathbb{Z},$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R},$ and let then $V_{j}\in B(H)$ be given by
-$V_{j}(f)=w_{j}\cdot(f\circ\alpha_{j})$
-for all $f\in H.$ For each $n\in\mathbb{N},i\in\mathbb{Z}$, let $W_{n,i}\in B(H)$ be given by
-$W_{n,i}:=V_{i}V_{i-1}\cdots V_{i-n+1}.$
-If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [16, Example 3.3] one can show that the conditions of Proposition 3.1 are satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R})$ in this case.
-On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 3.1 will be satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R}).$
-Next, for $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}.$
-Let $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$.
-We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
-$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in C_{b}(\Omega)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|_{\infty}\leq M_{y}\ \forall j\}.$
-The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
-For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
-[…]
-Then it is easily checked that $\{\tilde{p}_{(K_{1},K_{2},J)}^{3}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}$ satisfies the condition $(\mathcal{P}.)$
-For each $\ell\in\{1,\dots,N\}$ and $k\in\mathbb{N}$, we let $\tilde{\Phi}_{k,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
-$\tilde{\Phi}_{k,\ell}\big(\{f_{j}\}_{j\in\mathbb{Z}}\big)=\{f_{j-k}\circ\alpha_{\ell}^{k}\,\}_{j\in\mathbb{Z}},$
-for all $\{f_{j}\}_{j\in\mathbb{Z}}\subset\mathcal{A}$. Let now $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, set
-$\Phi_{n,\ell}=\tilde{\Phi}_{r_{\ell}n,\ell}.$
-Clearly, the system of isometric isomorphisms $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is disjoint aperiodic with respect to
-$\{\tilde{p}_{(K_{1},K_{2},J)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}.$
-Also, the condition (R) is satisfied in this case.
-For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
-$(b_{n,\ell})_{j+r_{\ell}n}=\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)\quad\text{for all }j\in\mathbb{Z}.$
-Since for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ the function $w_{j}^{(\ell)}>0,$ it follows that the condition $(C)$ is satisfied in this case.
-Now, for each $K_{2}\subseteq\Omega$ compact, we can find some $v_{K_{2}}\in C_{0}(\Omega)$ with $0\leq v_{K_{2}}\leq 1$ such that ${v_{K_{2}}}_{|_{K_{2}}}=1.$ Recalling that the approximate unit for $\mathcal{A}$ in this case is indexed over all triples $(K_{1},K_{2},J)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }J\in\mathbb{N},$ for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, and every fixed triple $(K_{1},K_{2},J)$ we let $b_{n,\ell,(K_{1},K_{2},J)}^{-1}\in\mathcal{A}_{1}$ be given by
-$(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=\frac{v_{K_{2}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)}$
-for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ and $(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=0$ else.
-Then, for all $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ we let $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
-$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a)\quad\text{for all }a\in\mathcal{A}.$
-The next corollary can be deduced by some calculations from Theorem 2.1.
-Corollary 3.3.
-Under the above notation and assumptions, the following statements are equivalent:
-(1) The sequences of operators
-$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
-are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
-(2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
-for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
-Example 3.4.
-Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$.
-Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 3.3 are satisfied in this case.
+In addition, we let $C_{c}(\Omega)$ be the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subsets $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ be such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
 4.1. Applications to generalized weighted composition operators on spaces of operator-valued continuous functions
 In this subsection, we let $C_{0}(\Omega,\mathcal{C})$ denote the non-unital normed algebra of all continuous $\mathcal{C}$-valued functions on $\Omega$ vanishing at infinity and equipped with the norm
-$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|$
+$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|,$
 and $C_{b}(\Omega,B(H))$ be the unital normed algebra consisting of all uniformly bounded $B(H)$-valued continuous functions on $\Omega$ equipped with the same norm $\|\cdot\|_{\infty}$.
@@
 Lemma 4.1.
-The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\in\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
+The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
 Proof.
@@
 $\leq\|f(y)-f(x)\|+\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}+\|(I-P_{m})f(x)\|$
-By [21, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
+By [29, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
 The collection $\{O_{x}\}_{x\in K_{1}}$ is an open cover of $K_{1}$, hence there exist a finite subcover $\{O_{x_{1}},\ldots,O_{x_{M}}\}.$ Let $\tilde{m}:=\max\{m_{x_{1}},\ldots,m_{x_{M}}\}.$ Then, by the above we get
@@
 $\|\tilde{P}^{3}_{(K_{1},K_{2},\tilde{m})}f-f\|_{\infty}<\varepsilon,\text{ so }\tilde{P}^{3}_{(K_{1},K_{2},m)}f\in O.$
-Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ We will assume that for every compact subset $K$ of $X$ and every $S\in\mathcal{F}$ there exists some $F\in\mathcal{F}$ with $F\subseteq S$ such that
-$K\cap\alpha_{n,\ell}(K)=\varnothing\quad\text{and}\quad\alpha_{n,\ell}^{-1}(\alpha_{n,s}(K))\cap K=\varnothing$
-for all $n\in F$ and each distinct $s,\ell\in\{1,\dots,N\}.$ Also, we consider families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
+Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ and families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
 $\Phi_{n,l}:\mathcal{A}_{1}\rightarrow\mathcal{A}_{1}\text{ be given by }\Phi_{t,l}(f):=U_{n,l}^{*}(f\circ\alpha_{n,l})U_{n,l}$
@@
 $D_{m}e_{j}\in H_{0}\text{ for all }j\in\{-m,\dots m\}\quad\text{and}\quad D_{m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\dots m\}.$
-By some calculations similar to those in the proof of [16, Theorem 3.2] and [17, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 2.2 are satisfied. ∎
+By some calculations similar to those in the proof of [20, Theorem 3.2] and [21, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 3.2 are satisfied. ∎
 Example 4.3.
 […]
-3.3. Applications to generalized weighted shifts on function sequence spaces
-In this subsection, we assume that $\Omega=\mathbb{R}.$ and we let now $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [15], for $\tau\in C_{b}(\mathbb{R}),$ we put
-$\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
-For each $f\in\mathcal{C}_{\tau}$ we define
-$\|f\|_{\tau}:=\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}.$
-Then, $\mathcal{C}_{\tau}$ is a Banach algebra [14]. We will call this algebra Segal algebra corresponding to $\tau$.
-As observed in [15], for a homeomorphism $\alpha$ of $\mathbb{R},$ we consider corresponding Segal algebra $\mathcal{C}_{\tau\circ\alpha^{k}}$ for all $k\in\mathbb{Z}$.
-Clearly, $f\cdot(\tau\circ\alpha^{k})\in\mathcal{C}$ for all $k\in\mathbb{Z}$ whenever $f\in C_{0}(\mathbb{R})$ since $\tau\circ\alpha^{k}\in C_{b}(\mathbb{R})$ for all $k\in\mathbb{Z}.$ By definition, $\|f\|_{\infty}\leq\|f\|_{\tau\circ\alpha^{k}}$ whenever $f\in\mathcal{C}_{\tau\circ\alpha^{k}}.$ We recall now the following lemmas from [15].
+5. Dynamics on weighted spaces of Radon measures
+If $T_{\alpha,w}$ denotes the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega),$ then the adjoint $T_{\alpha,w}^{*}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm.
+It is straightforward to check that
+$T_{\alpha,w}^{*}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
+for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
+${T_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
+and
+${S_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
+Throughout this paper, $M_{+}(\Omega)$ will denote the set of all non-negative Radon measures on $\Omega.$ For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$ A weight on $\Omega$ with respect to $\alpha$ is a continuous positive function $\rho$ on $\Omega$ satisfying that
+$(\rho\circ\alpha)(t)\leq C\rho(t)\quad\text{and}\quad(\rho\circ\alpha^{-1})(t)\leq C\rho(t)$
+for all $t\in\Omega$ and some constant $C>0$.
+If $\mu\in M(\Omega)$ and $\rho\in L^{1}(|\mu|)$, we let $\mu^{(\rho)}$ be the Borel measure on $\Omega$ given by
+$\mu^{(\rho)}(E)=\int_{E}\rho\,d\mu$
+for every Borel subset $E$ of $\Omega$ and we set
+$M_{\rho}(\Omega):=\left\{\mu\in M(\Omega)\;\middle|\;\rho\in L^{1}(|\mu|)\text{ and }\mu^{(\rho)}\text{ is regular}\right\}.$
+Let $\|\cdot\|_{\rho}:M_{\rho}(\Omega)\to\mathbb{R}^{+}$ be given by
+$\|\mu\|_{\rho}=|\mu^{(\rho)}|(\Omega).$
 Lemma 5.1.
-[15, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
+$M_{\rho}(\Omega)$ is a subspace and $\|\cdot\|_{\rho}$ is a norm on $M_{\rho}(\Omega).$
+Proof.
+Let $\mu,\nu\in M_{\rho}(\Omega)$ and $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu.$ It is easy to check that $|\mu+\nu|\leq|\mu|+|\nu|.$ Now, since $\rho\in L^{1}(|\mu|)\cap L^{1}(|\nu|),$ we get that $\rho\in L^{1}(|\mu+\nu|).$ Next, we have that $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big),$ is the Jordan decomposition of $\mu^{(\rho)}.$
+Therefore, $\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)},$ $\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)}\leq|\mu^{(\rho)}|,$ hence, if $|\mu^{(\rho)}|=0,$ then $\mu_{j}^{(\rho)}=0$ for all $j\in\{1,\ldots,4\},$ but this means that $\mu_{j}=0$ for all $j\in\{1,\ldots,4\}$ since $\rho$ is a positive continuous function and $\mu_{j}$ is a Borell measure.
+Thus, $\mu=0$ in this case.
+∎
+Lemma 5.2.
+Let $K$ be a compact subset of $\Omega,$ $\delta>0$ and $\mu\in M_{\rho}(\Omega).$ If $|\mu^{(\rho)}|(K)<\delta,$ then $|\mu|(K)\leq\dfrac{2\delta}{m_{K}},$ where $m_{K}:=\inf_{x\in K}\rho(x)>0.$
+Proof.
+Let again $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu,$ then $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)})$ is the Jordan decomposition of $\mu^{(\rho)}.$ If $|\mu^{(\rho)}|(K)<\delta,$ then $(\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)})(K)<\delta$ and $(\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)})(K)<\delta.$ Since $\rho$ is positive and continuous, we must have that $m_{K}>0,$ hence, the above inequalities give $(\mu_{1}+\mu_{2})(K)<\dfrac{\delta}{m_{K}}$ and $(\mu_{3}+\mu_{4})(K)<\dfrac{\delta}{m_{K}}.$ Now, since $|\mu|\leq\sum_{j=1}^{4}\mu_{j},$ the statement follows.
+∎
+Remark 5.3.
+If $\mu\in M_{\rho}(\Omega)$ and $\delta>0,$ it is not hard to check that there exists a compact subset $K\subseteq\Omega$ and $\nu\in M(\Omega)$ such that
+$|\nu|(K^{c})=0\quad\text{and}\quad|\mu^{(\rho)}-\nu^{(\rho)}|(\Omega)<\delta.$
+Indeed, since $\mu^{(\rho)}$ is a regular Borel measure, there exists a compact subset $K\subseteq\Omega$ such that
+$|\mu^{(\rho)}|(K^{c})<\frac{\delta}{2},$
+hence we can let $\nu$ be the Borel measure on $\Omega$ given by
+$\nu(E):=\mu(E\cap K),$
+for every Borel subset $E\subseteq\Omega.$
+Lemma 5.4.
+The operators $T_{\alpha,w}^{*}$ and $S_{\alpha,w}^{*}$ are bounded linear self-mappings on $M_{\rho}(\Omega).$
+Proof.
+Let $\mu\in M_{\rho}(\Omega)$ and
+$\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$
+be the Jordan decomposition of $\mu.$ Then $\rho\in\bigcap_{j=1}^{4}L^{1}(\mu_{j}).$ By the change of variables, we obtain
+$\int\rho\,(w\circ\alpha^{-1})\,d(\mu_{j}\circ\alpha^{-1})=\int(\rho\circ\alpha)\,w\,d\mu_{j}\leq$
+$C\Big(\sup_{t\in\Omega}w(t)\Big)\int\rho\,d\mu_{j}<\infty,$
+for all $j\in\{1,2,3,4\},$ because
+$(\rho\circ\alpha^{-1})(t)\leq C\rho(t),\quad\text{for all }t\in\Omega,$
+by the assumption. Thus, $\rho\in L^{1}\big(T_{\alpha,w}^{*}(\mu_{j})\big)$ for all $j\in\{1,2,3,4\}.$ Since
+$T^{*}_{\alpha,w}(\mu)=T^{*}_{\alpha,w}(\mu_{1})-T^{*}_{\alpha,w}(\mu_{2})+i\Big(T^{*}_{\alpha,w}(\mu_{3})-T^{*}_{\alpha,w}(\mu_{4})\Big)$
+is the Jordan decomposition of $T^{*}_{\alpha,w}(\mu)$ by the arguments from the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, it follows that $\rho\in L^{1}\big(|T_{\alpha,w}^{*}(\mu)|\big).$
+Next, given $\epsilon>0$ and a Borel subset $E$ of $\Omega,$ choose some open subset $\mathcal{O}\subseteq E$ and a compact subset $K\subseteq\Omega$ such that $K$ and $\alpha^{-1}(E)$ are subsets of $\mathcal{O}$ and
+$|\mu^{(\rho)}|(\mathcal{O}\setminus\alpha^{-1}(E)),|\mu^{(\rho)}|(\mathcal{O}\setminus K)<\varepsilon.$
+Since $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big)$ is the Jordan decomposition of $\mu^{(\rho)}$, we have that $|\mu_{j}^{(\rho)}|\leq|\mu^{(\rho)}|$ for all $j\in\{1,\dots,4\}$, hence
+$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus\alpha^{-1}(E)\big)<\varepsilon$
+and
+$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus K\big)<\varepsilon$
+for all $j\in\{1,\dots,4\}$. By the previous inequality, we get that
+$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)>C\Big(\sup_{t\in\Omega}w(t)\Big)\int_{\mathcal{O}\setminus\alpha^{-1}(E)}\rho\,d\mu_{j}\geq\int_{\alpha(\mathcal{O})\setminus(E)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big),$
+and similarly,
+$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)\geq\int_{\alpha(\mathcal{O})\setminus\alpha(K)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big)$
+for all $j\in\{1,\dots,4\}$.
+Since
+$\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}=\big(T^{*}_{\alpha,w}(\mu_{1})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{2})\big)^{(\rho)}+i\Big(\big(T^{*}_{\alpha,w}(\mu_{3})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{4})\big)^{(\rho)}\Big)$
+is the Jordan decomposition of $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$, we deduce that
+$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus E)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)$
+and
+$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus\alpha(K)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big).$
+Now, $\alpha(\mathcal{O})$ is open and $\alpha(K)$ is compact because $\alpha$ is a homeomorphism, and, moreover, $E\subseteq\alpha(\mathcal{O}),$ hence $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$ is regular.
+Finally, by the above,
+$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\leq 4C\Big(\sup_{t\in\Omega}w(t)\Big)|\mu^{(\rho)}|.$
+Theorem 5.5.
+Under the above notation and assumptions, the following statements are equivalent.
   1. (1)
-    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}=1$ on $K_{\epsilon_{2}}^{(\tau)}$,
+    The set of periodic elements of $T_{\alpha,\omega}^{*}$ is dense in $M_{\rho}(\Omega)$.
   2. (2)
-    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in C_{c}(\mathbb{R}),$
-  3. (3)
-    ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\subseteq|\tau|^{-1}(\left[0,\epsilon_{1}\right])).$
-Lemma 3.9.
-[15, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
-Lemma 3.10.
-[15, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$.
-Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
-As in [15], let
-${\rm c}_{0}^{\mathcal{C},\tau}:=\Big\{s:=(s_{k})\in\prod_{k\in\mathbb{Z}}\mathcal{C}_{\tau\circ\alpha^{k}}:\,\lim_{k\rightarrow\infty}\|s_{k}\|_{\tau\circ\alpha^{k}}=\lim_{k\rightarrow\infty}\|s_{-k}\|_{\tau\circ\alpha^{-k}}=0\Big\}$
-and we equip ${\rm c_{0}}^{\mathcal{C},\tau}$ with the norm
-$\|s\|_{0}:=\sup_{k\in\mathbb{Z}}\|s_{k}\|_{\tau\circ\alpha^{k}}.$
-Since by [14] we have that $\mathcal{C}_{\tau\circ\alpha^{k}}$ is Banach algebra for all $k\in\mathbb{Z},$ it is not hard to check that ${\rm c}_{0}^{\mathcal{C},\tau}$ is a non-unital Banach algebra.
-We will from now on denote ${\rm c_{0}}^{\mathcal{C},\tau}$ by $\mathcal{A}.$ Moreover, we let $\mathcal{A}_{1}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is straightforward to check that the condition (E) is satisfied also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}$ (and $\mathcal{A}_{1}$ is as above in the previous case).
-For $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1},J\in\mathbb{N},$ and any collection of compact subsets $K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}$ of $\mathbb{R},$ we let $\tilde{p}_{(K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1}}\in\mathcal{A}$ be given by
+    For every compact subset $K$ of $\Omega$ and every measure $\mu\in M_{\rho}(\Omega)$ with $|\mu|(K^{c})=0$, there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences $\{A_{k}\}_{k\in\mathbb{N}},\{B_{k}\}_{k\in\mathbb{N}}$ of Borel subsets of $K$ such that $\alpha^{n_{k}}(K)\cap K=\varnothing$ for all $k\in\mathbb{N}$ and
+    $\lim_{k\to\infty}|\mu|(B_{k})=\lim_{k\to\infty}|\mu|(A_{k})=0,$
+    and
+    $\lim_{k\to\infty}\sup_{t\in A_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\left(\prod_{r=0}^{\ell n_{k}-1}(w\circ{\alpha^{r}})(t)\right)=$
+    $\lim_{k\to\infty}\sup_{t\in B_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})(t)\left(\prod_{r=1}^{\ell n_{k}}(w\circ{\alpha^{-r}})\right)^{-1}(t)=0.$
+Proof.
+First we will prove $1.\Rightarrow 2.$ If 1.
+holds, that is if the set of periodic elements of $T_{\alpha,\omega}^{*}$ is dense in $M_{\rho}(\Omega)$, then given a compact subset $K\subseteq\Omega$ and $\mu\in M_{\rho}(\Omega)$ with $|\mu|(K^{c})=0$, for each $k\in\mathbb{N}$ we can find some $\gamma_{k}\in M_{\rho}(\Omega)$ with
+$T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\gamma_{k}\quad\text{for all }\ell\in\mathbb{N}\quad\text{ and some }n_{k}\in\mathbb{N},$
+and such that
+$|\gamma_{k}^{(\rho)}-\mu^{(\rho)}|(\Omega)<\frac{1}{4^{k}}.$
+Then
+$T_{\alpha,\omega}^{*\,\ell(mn_{k})}(\gamma_{k})=\gamma_{k}\quad\text{for all }\ell,m\in\mathbb{N},$
+hence the sequence $\{n_{k}\}_{1}$ can be chosen to be strictly increasing, and such that
+$\alpha^{r}(K)\cap K=\varnothing$
+for all $r\geq n_{k}$ due to the fact that $\alpha$ is aperiodic. Notice that this gives that
+$\alpha^{j_{1}n_{k}}(K)\cap\alpha^{j_{2}n_{k}}(K)=\varnothing$
+for each distinct $j_{1},j_{2}\in\mathbb{Z}$ and all $k\in\mathbb{N}$. Moreover, since $|\mu|(K^{c})=0$, then
+$|\mu^{(p)}|(K^{c})=0,$
+hence
+$|\gamma_{k}^{(p)}|(K^{c})<\frac{1}{4^{k}}\quad\text{for all }k\in\mathbb{N}.$
+Taking all this into account, for each $k\in\mathbb{N}$ we then get
+$|\gamma_{k}^{(\rho)}|(K^{c})\geq\sum_{\ell=1}^{\infty}|\gamma_{k}^{(\rho)}|\big(\alpha^{\ell n_{k}}(K)\big)+\sum_{\ell=1}^{\infty}|\gamma_{k}^{(\rho)}|\big(\alpha^{-\ell n_{k}}(K)\big)$
+$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\big(\alpha^{\ell n_{k}}(K)\big)+\sum_{\ell=1}^{\infty}\left|\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\big(\alpha^{-\ell n_{k}}(K)\big).$
+For each $k,\ell\in\mathbb{N}$ and $j\in\{1,2\}$, let $\eta_{j,\pm}^{(k,\ell)}\in M(\Omega)$ be given by
+$\eta_{j,\pm}^{(k,\ell)}(E)=\int_{E}\prod_{r=1}^{\ell n_{k}}\big(w\circ\alpha^{\ell n_{k}-r}\big)^{-1}\,d\big(\gamma_{j,\pm}^{(k)}\circ\alpha^{\ell n_{k}}\big)$
+for every Borel subset $E$ of $\Omega$, where
+$\gamma_{k}=\gamma_{1,+}^{(k)}-\gamma_{1,-}^{(k)}+i\big(\gamma_{2,+}^{(k)}-\gamma_{2,-}^{(k)}\big)$
+is the Jordan decomposition of $\gamma_{k}$.
+Then, by the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we have that
+$S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\eta_{1,+}^{(k,\ell)}-\eta_{1,-}^{(k,\ell)}+i\big(\eta_{2,+}^{(k,\ell)}-\eta_{2,-}^{(k,\ell)}\big)$
+is the Jordan decomposition of $S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})$.
+Hence,
+$\big(\eta_{1,+}^{(k,\ell)}\big)^{(\rho)}-\big(\eta_{1,-}^{(k,\ell)}\big)^{(\rho)}+i\left[\big(\eta_{2,+}^{(k,\ell)}\big)^{(\rho)}-\big(\eta_{2,-}^{(k,\ell)}\big)^{(\rho)}\right]$
+is the Jordan decomposition of $\big(S_{2,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}$. Therefore,
+$\big(\eta_{j,+}^{(k,\ell)}\big)^{(\rho)}+\big(\eta_{j,-}^{(k,\ell)}\big)^{(\rho)}\leq\left|\big(S_{2,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\quad\text{for }j\in\{1,2\}.$
+Hence,
+$\sum_{\ell=1}^{\infty}\Big((\eta_{j,+}^{(k,\ell)})^{(\rho)}+(\eta_{j,-}^{(k,\ell)})^{(\rho)}\Big)\big(\alpha^{-\ell n_{k}}(K)\big)<\frac{1}{4^{k}}$
+for all $k\in\mathbb{N}$ and $j\in\{1,2\}$. Thus,
+$\frac{1}{4^{k}}>\sum_{\ell=1}^{\infty}\int_{\alpha^{-\ell n_{k}}(K)}\rho\,\Big(\prod_{r=1}^{\ell n_{k}}\big(w\circ\alpha^{\ell n_{k}-r}\big)\Big)^{-1}\,d\big((\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)})\circ\alpha^{\ell n_{k}}\big)=$
+$\sum_{\ell=1}^{\infty}\int_{K}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}\,d\big(\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)}\big)=$
+$\int_{K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}\,d\big(\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)}\big)$
+for each $k\in\mathbb{N}$ and $j\in\{1,2\}$, where in the first equality we have applied the change of variables and in the second equality we have applied the Monotone Convergence Theorem.
+For each $k\in\mathbb{N}$, let
+$B_{k}=\left\{t\in K\ \middle|\ \sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})(t)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}(t)>\frac{1}{2^{k}}\right\}.$
+Since
+$|\gamma_{k}^{(\rho)}-\mu^{(\rho)}|(K)<\frac{1}{4^{k}},$
+by Lemma 5.2 we have
+$|\gamma_{k}-\mu|(K)<\frac{2}{m_{K}4^{k}},$
+hence, by similar arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we can deduce that
+$|\mu|(B_{k})<\frac{1}{2^{k-1}}+\frac{2}{m_{K}4^{k}}\quad\text{for all }k\in\mathbb{N}.$
+Next, by observing that
+$T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\Theta_{1,+}^{(k,\ell)}-\Theta_{1,-}^{(k,\ell)}+i\big(\Theta_{2,+}^{(k,\ell)}-\Theta_{2,-}^{(k,\ell)}\big)$
+is the Jordan decomposition of $T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})$, where
+$\Theta_{j,\pm}^{(k,\ell)}(E)=\int_{E}\prod_{r=0}^{\ell n_{k}-1}\big(w\circ\alpha^{r-\ell n_{k}}\big)\,d\big(\gamma_{j,\pm}^{(k)}\circ\alpha^{-\ell n_{k}}\big)$
+for each $\ell,k\in\mathbb{N}$, $j\in\{1,2\}$ and every Borel subset $E$ of $\Omega$, and using that
+$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(p)}\right|\big(\alpha^{\ell n_{k}}(K)\big)<\frac{1}{4^{k}},$
+in a similar way as above we obtain that
+$|\mu|(A_{k})<\frac{1}{2^{k-1}}+\frac{2}{m_{K}4^{k}},$
+where
+$A_{k}:=\left\{t\in K\ \middle|\ \sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\prod_{r=0}^{\ell n_{k}-1}\big(w\circ\alpha^{\,r}\big)(t)>\frac{1}{2^{k}}\right\}$
+for all $k\in\mathbb{N}$. This proves $1.\Rightarrow 2.$
+Next we will prove $2.\Rightarrow 1.$ Suppose that 2.
+holds.
+Given an open non-empty subset $O$ of $M_{\rho}(\Omega)$, by Remark 5.3 there exists a compact subset $K\subseteq\Omega$ and $\mu\in O$ such that $|\mu|(K^{c})=0$.
+Choose a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of Borel subsets $\{A_{n}\}_{n},\{B_{k}\}_{k}$ of $K$ satisfying the assumption in (2) with respect to $\mu$ and $K$.
+Let
+$C_{k}=A_{k}\cup B_{k}\quad\text{for each }k\in\mathbb{N},$
+and let $\mu_{k}\in M_{\rho}(\Omega)$ be given by
+$\mu_{k}(E)=\mu(E\cap C_{k}^{c})$
+for every Borel subset $E$ of $\Omega$. Then, $|\mu_{k}|(K^{c}\cup C_{k})=0$ for each $k\in\mathbb{N}$.
+Let
+$\mu_{k}=\mu_{1,+}^{(k)}-\mu_{1,-}^{(k)}+i\big(\mu_{2,+}^{(k)}-\mu_{2,-}^{(k)}\big)$
+be the Jordan decomposition of $\mu_{k}$ for each $k\in\mathbb{N}$. As above in the proof of the implication $(1)\Rightarrow(2)$, if
+$\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}=\widetilde{\Theta}_{1,+}^{(k,\ell)}-\widetilde{\Theta}_{1,-}^{(k,\ell)}+i\big(\widetilde{\Theta}_{2,+}^{(k,\ell)}-\widetilde{\Theta}_{2,-}^{(k,\ell)}\big)$
+and
+$\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}=\widetilde{\eta}_{1,+}^{(k,\ell)}-\widetilde{\eta}_{1,-}^{(k,\ell)}+i\big(\widetilde{\eta}_{2,+}^{(k,\ell)}-\widetilde{\eta}_{2,-}^{(k,\ell)}\big),$
+are the Jordan decompositions of $\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}$ and $\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}$, respectively, for each $\ell,k\in\mathbb{N}$, then
+$\widetilde{\Theta}_{j,\pm}^{(k,\ell)}(E)=\int_{E}\rho\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{\,r-n_{k}})\Big)\,d\big(\mu_{j,\pm}^{(k)}\circ\alpha^{-\ell n_{k}}\big)$
+and
+$\widetilde{\eta}_{j,\pm}^{(k,\ell)}(E)=\int_{E}\rho\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{\ell n_{k}-r})^{-1}\Big)\,d\big(\mu_{j,\pm}^{(k)}\circ\alpha^{\ell n_{k}}\big)$
+for each $k,\ell\in\mathbb{N},j\in\{1,2\}$ and every Borel subset $E$ of $\Omega$. Therefore, for each $k\in\mathbb{N}$,
+$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}\right|(\Omega)+\sum_{\ell=1}^{\infty}\left|\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}\right|(\Omega)$
+$\leq\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\big(\widetilde{\Theta}_{j,+}^{(k,\ell)}+\widetilde{\Theta}_{j,-}^{(k,\ell)}\big)(\Omega)+\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\big(\widetilde{\eta}_{j,+}^{(k,\ell)}+\widetilde{\eta}_{j,-}^{(k,\ell)}\big)(\Omega)$
+$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{\Omega}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{\Omega}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{C_{k}^{c}\cap K}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{C_{k}^{c}\cap K}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\sum_{j=1}^{2}\int_{C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\sum_{j=1}^{2}\int_{C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
+$\leq\sum_{j=1}^{2}\sup_{t\in C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})(t)\Big)\,(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)$
+$\sum_{j=1}^{2}\sup_{t\in C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{-\ell n_{k}})(t)\Big)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}(t)\Big)\,(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)$
+$\leq 2\sup_{t\in A_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{\ell n_{k}})(t)\Big)\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})(t)\Big)\,|\mu|(\Omega)$
+$2\sup_{t\in B_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{-\ell n_{k}})(t)\Big)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}(t)\Big)\,|\mu|(\Omega)\longrightarrow 0,\quad k\to\infty,$
+where in the first equality we have applied the change of variables, in the second equality we have used the fact that $|\mu_{k}|(K^{c}\cup C_{k})=0$ for each $k\in\mathbb{N}$, in the third equality we have applied the Monotone Convergence Theorem, and in the last inequality we have used the fact that
+$(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)\leq|\mu^{(k)}|(\Omega)\leq|\mu|(\Omega)$
+for all $k\in\mathbb{N}$ and $j\in\{1,2\}$.
+Further, if
+$\mu=\mu_{1,+}-\mu_{1,-}+i(\mu_{2,+}-\mu_{2,-})$
+is the Jordan decomposition of $\mu$, and $\widetilde{\mu}_{k,j,\pm}^{(p)}$ is the Borel measure on $\Omega$ given by
+$\widetilde{\mu}_{k,j,\pm}^{(\rho)}(E)=\int_{E\cap C_{k}}\rho\,d\mu_{j,\pm},$
+for every Borel subset $E$ of $\Omega,$ $k\in\mathbb{N}$ and $j\in\{1,2\}$, then, clearly,
+$\widetilde{\mu}_{k,1,+}^{(\rho)}-\widetilde{\mu}_{k,1,-}^{(\rho)}+i\big(\widetilde{\mu}_{k,2,+}^{(\rho)}-\widetilde{\mu}_{k,2,-}^{(\rho)}\big)$
+is the Jordan decomposition of $\mu^{(\rho)}-\mu_{k}^{(\rho)}$ for each $k\in\mathbb{N}$.
+Since
+$\int_{C_{k}}\rho\,d\mu_{j,\pm}\leq\sup_{t\in K}\rho(t)\,\mu_{j,\pm}(C_{k})\leq\sup_{t\in K}\rho(t)\,|\mu|(C_{k})\longrightarrow 0,\quad k\to\infty,$
+we deduce that
+$\big\|\mu^{(p)}-\mu_{k}^{(p)}\big\|(\Omega)\longrightarrow 0,\quad k\to\infty.$
+As in the proof of [23, Proposition 10.3], for each $k\in\mathbb{N},$ set
+$v_{k}=\mu_{k}+\sum_{l=1}^{\infty}T_{\alpha,w}^{*^{ln_{k}}}(\mu_{k})+\sum_{l=1}^{\infty}S_{\alpha,w}^{*^{ln_{k}}}(\mu_{k}).$
+It is straightforward to check that $v_{k}=T_{\alpha,w}^{*ln_{k}}v_{k}$ for all $l,k\in\mathbb{N}.$ Moreover,
+$v_{k}\longrightarrow\mu,\quad k\to\infty.$
+This proves $2.\Rightarrow 1.$
+Let $N\in\mathbb{N}$ and $\{\alpha_{z,1}\}_{z\in\mathbb{Z}},\dots,\{\alpha_{z,N}\}_{z\in\mathbb{Z}}$ be families of homeomorphisms of $\Omega$ which are $d\mathcal{F}$-aperiodic, that is for each compact subset $K\subseteq\Omega$ and every $E\in\mathcal{F}$, there exists some $F\subseteq E$ with $F\in\mathcal{F}$ such that
+$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing$
+for every $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$ and in addition
+$\alpha_{z,\ell}(K)\cap K=\varnothing$
+for all $z\in F$ and $\ell\in\{1,\dots,N\}$.
+We will assume that $\rho$ is a weight with respect to $\alpha_{z,\ell}$ for all $z\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$.
+Let $\{w_{z,1}\}_{z\in\mathbb{Z}},\dots,\{w_{z,N}\}_{z\in\mathbb{Z}}$ be families of continuous bounded positive functions on $\Omega$ such that $w_{z,\ell}^{-1}$ is also bounded for each $z\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$.
+We let
+$T^{*}_{z,\ell}:=T^{*}_{\alpha_{z,\ell},\,w_{z,\ell}}\quad\text{for all }z\in\mathbb{Z}\text{ and }\ell\in\{1,\dots,N\}.$
+Theorem 5.6.
+Under the above notation and assumptions, the following are equivalent.
+(1) The families $\{T^{*}_{z,1}\}_{z\in\mathbb{Z}},\dots,\{T^{*}_{z,N}\}_{z\in\mathbb{Z}}$ are $d\mathcal{F}$-semi-transitive on $M_{\rho}(\Omega)$.
+(2) For each compact subset $K\subseteq\Omega,\varepsilon>0$ and any measures $\mu,\nu_{1},\dots,\nu_{N}\in M_{\rho}(\Omega)$ with
+$|\mu|(K^{c})=|\nu_{1}|(K^{c})=\dots=|\nu_{N}|(K^{c})=0,$
+there exists some $F\in\mathcal{F}$ and families of Borel subsets of $K$
+$\{A_{z}\}_{z\in F},\quad\{B_{z,1}\}_{z\in F},\dots,\{B_{z,N}\}_{z\in F}$
+such that
+$|\mu|(A_{z})<2N\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right),$
+$|\nu_{s}|(B_{z,s})<2(N+1)\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right),$
+and
+$\left(\sup_{t\in A_{z}^{c}\cap K}\big((\rho\circ\alpha_{z,\ell})(t)\,w_{z,\ell}(t)\big)\right)\left(\sup_{t\in B_{z,s}^{c}\cap K}\big((\rho\circ\alpha^{-1}_{z,s})(t)\,(w_{z,s}\circ\alpha^{-1}_{z,s})(t)\big)\right)<\varepsilon$
+for all $z\in F$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for all $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$, it holds that
+$\sup_{t\in B_{z,s}^{c}\cap K}\left((\rho\circ\alpha_{z,\ell}\circ\alpha^{-1}_{z,s})(t)\,\frac{\big(w_{z,\ell}\circ\alpha^{-1}_{z,s}\big)(t)}{\big(w_{z,s}\circ\alpha^{-1}_{z,s}\big)(t)}\right)<\sqrt{\varepsilon}.$
+Proof.
+First we will prove $(1)\Rightarrow(2)$. Given $\varepsilon>0,$ a compact subset $K\subseteq\Omega$ and measures $\mu,\nu_{1},\dots,\nu_{N}\in M(\Omega)$ such that
+$0=|\mu|(K^{c})=|\nu_{1}|(K^{c})=\dots=|\nu_{N}|(K^{c}),$
+we can find some $E\in\mathcal{F}$ and some $\lambda_{z}\in\mathbb{R}^{+}$ such that
+$\frac{1}{\lambda_{z}}S^{*}_{z,1}\big(B(\nu_{1},\varepsilon)\big)\cap\dots\cap\frac{1}{\lambda_{z}}S^{*}_{z,N}\big(B(\nu_{N},\varepsilon)\big)\cap B(\mu,\varepsilon)\neq\varnothing$
+for all $z\in E$, where $B(\nu_{\ell},\varepsilon)$ is the open ball in $M_{\rho}(\Omega)$ with centre in $\nu_{\ell}$ and radius $\varepsilon$ for all $\ell\in\{1,\dots,N\}$.
+Further, since $\{\alpha_{z,1}\}_{z\in\mathbb{R}},\dots,\{\alpha_{z,N}\}_{z\in\mathbb{R}}$ is a $d\mathcal{F}$-aperiodic system, we can find some $F\subseteq E$ with $F\in\mathcal{F}$ such that
+$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing$
+for every $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$ and in addition
+$\alpha_{z,\ell}(K)\cap K=\varnothing$
+for all $z\in F$ and $\ell\in\{1,\dots,N\}$. Next, for each $z\in F$ there exists some $\eta_{z}\in M(\Omega)$ such that
+$\big|\eta_{z}^{(\rho)}-\mu^{(\rho)}\big|(\Omega)<\varepsilon\quad\text{and}\quad\big|\lambda_{z}\gamma_{z,\ell}^{(\rho)}-\nu_{\ell}^{(\rho)}\big|(\Omega)<\varepsilon\quad\text{for all }\ell\in\{1,\dots,N\},$
+where
+$\gamma_{z,\ell}^{(\rho)}=\big(T^{*}_{z,\ell}(\eta_{z})\big)^{(\rho)},$
+that is
+$\gamma_{z,\ell}^{(\rho)}(E)=\int_{E}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\,d\big(\eta_{z}\circ\alpha_{z,\ell}^{-1}\big)$
+for every measurable subset $E$ of $\Omega$.
+Observe that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $z\in F$ we have
+$T^{*}_{z,\ell}\Big(S^{*}_{z,s}\big(B(\nu_{s},\varepsilon)\big)\Big)\cap B(\nu_{\ell},\varepsilon)\neq\varnothing,$
+since
+$S^{*}_{z,s}\big(B(\nu_{s},\varepsilon)\big)\cap S^{*}_{z,\ell}\big(B(\nu_{\ell},\varepsilon)\big)\neq\varnothing.$
+Hence, for each $z\in F$, we can find some $\beta_{z,s}\in B(\nu_{s},\varepsilon)$ with
+$\Big|\big(T^{*}_{z,\ell}(S^{*}_{z,s}(\beta_{z,s}))\big)^{(\rho)}-\nu_{\ell}^{(\rho)}\Big|(\Omega)<\varepsilon.$
+Let
+$\Theta_{z,\ell,s}:=T^{*}_{z,\ell}\big(S^{*}_{z,s}(\beta_{z,s})\big).$
+We notice that for each Borel subset $E$ of $\Omega$ we have
+$\Theta_{z,\ell,s}^{(\rho)}(E)=\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\rho\;d\Big(S^{*}_{z,s}(\beta_{z,s})\circ\alpha_{z,\ell}^{-1}\Big)=$
+$\int_{\alpha_{z,\ell}^{-1}(E)}(\rho\circ\alpha_{z,\ell})\,w_{z,\ell}\;d\Big(S^{*}_{z,s}(\beta_{z,s})\Big)=\int_{\alpha_{z,\ell}^{-1}(E)}(\rho\circ\alpha_{z,\ell})\,w_{z,\ell}\,w_{z,s}^{-1}\;d\big(\beta_{z,s}\circ\alpha_{z,s}\big)=$
+$\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,\rho\;d\big(\beta_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big).$
+By using that $\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}$ is a homeomorphism of $\Omega$ with
+$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing,$
+together with the fact that
+$|\beta_{z,s}^{(\rho)}-\nu_{s}^{(\rho)}|(K^{c})=|\beta_{z,s}^{(\rho)}|(K^{c})<\varepsilon,$
+by exactly the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we get that
+$\Theta_{z,\ell,s,j,\pm}^{(\rho)}(E)=\int_{E}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,d\big(\beta_{z,s,j,\pm}\circ\alpha_{z,s}^{-1}\circ\alpha_{z,\ell}^{-1}\big)$
+for $j\in\{1,2\}$, where
+$\Theta_{z,\ell,s}^{(\rho)}=\Theta_{z,\ell,s,1,+}^{(\rho)}-\Theta_{z,\ell,s,1,-}^{(\rho)}+i(\Theta_{z,\ell,s,2,+}^{(\rho)}-\Theta_{z,\ell,s,2,-}^{(\rho)})$
+and
+$\beta_{z,s}=\beta_{z,s,1,+}-\beta_{z,s,1,-}+i(\beta_{z,s,2,+}-\beta_{z,s,2,-}),$
+are Jordan decompositions of $\Theta_{z,\ell,s}^{(\rho)}$ and $\beta_{z,s},$ respectively. Since
+$|\Theta_{z,\ell,s}^{(\rho)}|(K^{c})=\big|\Theta_{z,\ell,s}^{(\rho)}-\nu_{\ell}^{(\rho)}\big|(K^{c})<\varepsilon$
+and
+$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing,$
+we must have that
+$\varepsilon>\int_{\alpha_{z,\ell}(\alpha_{z,s}^{-1}(K))}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,d\big((\beta_{z,s,j,+}+\beta_{z,s,j,-})\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)=$
+$\int_{K}(\rho\circ\alpha_{z},l\circ\alpha_{z,s}^{-1})\big(w_{z,\ell}\circ\alpha_{z,s}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,s}^{-1}\big)^{-1}\,d\big(\beta_{z,s,j,+}+\beta_{z,s,j,-})$
+for $j\in\{1,2\}.$ Let
+$C_{z,\ell,s}:=\left\{t\in K\;\middle|\;(\rho\circ\alpha_{z,l}\circ\alpha_{z,s}^{-1})(t)\frac{\big(w_{z,\ell}\circ\alpha_{z,s}^{-1}\big)(t)}{\big(w_{z,s}\circ\alpha_{z,s}^{-1}\big)(t)}>\sqrt{\varepsilon}\right\}.$
+Then
+$\big(\beta_{z,s,j,+}+\beta_{z,s,j,-}\big)(C_{z,\ell,s})<\sqrt{\varepsilon}\quad\text{for }j\in\{1,2\}.$
+This gives that
+$|\beta_{z,s}|(C_{z,\ell,s})<2\sqrt{\varepsilon}.$
+Now, since
+$C_{z,\ell,s}\subseteq K,$
+we get
+$|\nu_{s}|(C_{z,\ell,s})\leq|\nu_{s}-\beta_{z,s}|(C_{z,\ell,s})+|\beta_{z,s}|(C_{z,\ell,s})\leq$
+$\leq|v_{s}-\beta_{z,s}|(K)+2\sqrt{\varepsilon}\leq\dfrac{2\varepsilon}{m_{K}}+2\sqrt{\varepsilon},$
+because
+$|v_{s}^{(\rho)}-\beta_{z,s}^{(\rho)}|(K)\leq|v_{s}^{(\rho)}-\beta_{z,s}^{(\rho)}|(\Omega)<\varepsilon,$
+and by Lemma 5.2 we then have that
+$|v_{s}-\beta_{z,s}|(K)\leq\dfrac{2\varepsilon}{m_{K}}.$
+By exactly the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$ and [24, Proposition 4.1] part $i)\Rightarrow ii)$ , we can deduce that
+$|\eta_{z}|\big(\widetilde{A}_{z,\ell}\big)<2\sqrt{\varepsilon}\quad\text{and}\quad|\lambda_{z}\gamma_{z,\ell}|\big(D_{z,\ell}\big)<2\sqrt{\varepsilon},$
+where
+$\widetilde{A}_{z,\ell}:=\left\{t\in K\;\middle|\;\big((\rho\circ\alpha_{z,\ell})(t)\big)\big(w_{z,\ell}(t)\big)>\frac{\sqrt{\varepsilon}}{\lambda_{z}}\right\}$
+and
+$D_{z,s}:=\left\{t\in K\;\middle|\;\big((\rho\circ\alpha_{z,s}^{-1})(t)\big)\big((w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\big)>\lambda_{z}\sqrt{\varepsilon}\right\}$
+for all $z\in F$ and $\ell,s\in\{1,\dots,N\}$. Again, since
+$|\eta_{z}^{(\rho)}-\mu^{(\rho)}|(\Omega)<\varepsilon\quad\text{and}\quad|\lambda_{z}\gamma_{z,\ell}^{(\rho)}-\nu_{\ell}^{(\rho)}|(\Omega)<\varepsilon,$
+we can conclude that
+$|\mu|\big(\widetilde{A}_{z,\ell}\big)<2\sqrt{\varepsilon}+\frac{2\varepsilon}{m_{K}}$
+and
+$|\nu_{\ell}|\big(D_{z,\ell}\big)<2\sqrt{\varepsilon}+\frac{2\varepsilon}{m_{K}}$
+for all $z\in F$ and $\ell\in\{1,\dots,N\}$.
+Set
+$A_{z}:=\bigcup_{\ell=1}^{N}\widetilde{A}_{z,\ell}$
+and
+$B_{z,s}:=D_{z,s}\;\cup\;\big(\bigcup_{\ell=1}^{N}C_{z,\ell,s}\big),$
+for each $z\in F$ and $s\in\{1,\dots,N\}.$ Then
+$|\mu|(A_{z})<2N\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right)$
+and
+$|\nu_{s}|(B_{z,s})<2(N+1)\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right)$
 […]
-(where $\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is the function from Lemma 3.8 with compact ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$).
-By the proof of Lemma 3.9 (see [15]) one can deduce that the condition ($\mathcal{P}$) is satisfied in this case.
-Let now $\alpha=\alpha_{1}=\cdots=\alpha_{N}.$ For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\Phi_{n,\ell}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is not hard to check that the system $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is again a system of isometric isomorphisms also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}.$ Moreover, it is disjoint apreiodic and satisfies the condition (R) with respect to
-$\{\tilde{p}_{((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}\}.$
-For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\mathbb{R}),b_{n,\ell}\in\mathcal{A}_{1}$ and $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be as above in the previous case (when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$).
-Now, given $\epsilon_{1}\in(0,1),$ we choose some $\epsilon_{3}\in(\epsilon_{1},1).$ Since ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$ and ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is compact, by Lemma 3.8, we can find some function $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\in\mathcal{C}_{\tau\circ\alpha^{j}}$ such that $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}(t)=1$ for all $t\in supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}},\text{ }supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}$ is compact and $supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\subseteq|\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{3}\right]).$ Therefore, for each $n\in\mathbb{N},\ell\in\{1,\dots,N\},J\in\mathbb{N},\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}$ and any collection of compact subsets
-$K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})},$
-we let $b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1}\in\mathcal{A}_{1}$ be given by
-$(b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1})_{j+r_{\ell}n}=\frac{v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j+r_{\ell}n})},\epsilon_{1}}\bigr),\epsilon_{3}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha^{n-k}\big)}$
-for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ (where we choose $\epsilon_{3}\in(\epsilon_{1},1)$) and $(b_{n,\ell,(\epsilon_{2},\epsilon_{1}J)}^{-1})_{j+r_{\ell}n}=0$ else.
-By some calculations, it is not hard to deduce the next corollary as a corollary of Theorem 2.1.
-Corollary 3.11.
-Under the above notation and assumptions, the following statements are equivalent:
-(1) The sequences of operators
-$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
-are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
-(2) For every $J\in\mathbb{N}$ and finite collection
-$\{K_{\epsilon_{J}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{-J+1}}^{(\tau\circ\alpha^{-J+1})},\ldots,K_{\epsilon_{0}}^{\tau},\ldots,K_{\epsilon_{J}}^{(\tau\circ\alpha^{J})}\}$
-of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
-for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
-$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$
+Now we will prove $(2)\Rightarrow(1)$.
+Given non-empty open subsets $O,O_{1},\dots,O_{N}$ of $M(\Omega)$, by Remark 5.3 we can find some $\mu\in O,$ $\nu_{\ell}\in O_{\ell}(\ell\in\{1,\dots,N\})$, and a compact $K\subseteq O$ such that
+$|\mu|(K^{c})=0\quad\text{and}\quad|\nu_{\ell}|(K^{c})=0\quad\text{for each }\ell\in\{1,\dots,N\}.$
+Also, there exists some $\delta>0$ such that
+$B(\mu,\delta)\subseteq O\quad\text{and}\quad B(\nu_{\ell},\delta)\subseteq O_{\ell}$
+for each $\ell\in\{1,\dots,N\}$. Let
+$\varepsilon:=\frac{\delta^{2}}{\left(2(N+1)\left(2+\frac{1}{m_{K}}\right)+\delta\right)^{2}}.$
+Choose $F\in\mathcal{F}$ and the families $\{A_{z}\}_{z\in F}$, $\{B_{z,1}\}_{z\in F},\dots,\{B_{z,N}\}_{z\in F}$ satisfying the assumptions of (2) with respect to $K$ and $\varepsilon$.
+Let $\widetilde{\mu}_{z},\widetilde{\nu}_{z,1},\dots,\widetilde{\nu}_{z,N}$ be the measures in $M(\Omega)$ given by
+$\widetilde{\mu}_{z}(E)=\mu(E\cap A_{z}^{c}\cap K)$
+and
+$\widetilde{\nu}_{z,\ell}(E)=\nu_{\ell}(E\cap B_{z,\ell}^{c}\cap K)$
+for every Borel subset $E$ of $\Omega$, and each $z\in F,\ell\in\{1,\dots,N\}$.
+Then, for each distinct $s,\ell\in\{1,\dots,N\}$,
+$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\rho\,d\big(\widetilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)$
+for all $z\in F$ and every Borel subset $E$ of $\Omega$. Hence,
+$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E\cap\alpha_{z,\ell}\big(\alpha_{z,s}^{-1}(B_{z,s}^{c}\cap K)\big)}\rho\frac{(w_{z,\ell}\circ\alpha_{z,\ell}^{-1})}{(w_{z,s}\circ\alpha_{z,\ell}^{-1})}\,d\big(\tilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)=$
+$\int_{\alpha_{z,s}\!\big(\alpha_{z,\ell}^{-1}(E)\big)\cap\big(B_{z,s}^{c}\cap K\big)}(\rho\circ\alpha_{z,l}\circ\alpha_{z,s}^{-1})\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})}{(w_{z,s}\circ\alpha_{z,s}^{-1})}\,d\tilde{v}_{z,s}$
+for every $z\in F,$ each distinct $s,\ell\in\{1,\dots,N\}$ and every Borel subset $E$ of $\Omega$.
+By exactly the same arguments as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ we can conclude that
+$\Big\|\big(T^{*}_{z,\ell}(S^{*}_{z,s}(\widetilde{\nu}_{z,s}))\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\left((\rho\circ\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(t)\right)\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})(t)}{(w_{z,s}\circ\alpha_{z,s}^{-1})(t)}\|\nu_{s}\|.$
+Similarly, as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ and [24, Proposition 4.1] part $ii)\Rightarrow i)$, we can deduce that
+$\Big\|\big(T^{*}_{z,\ell}(\widetilde{\mu}_{z})\big)^{(\rho)}\Big\|\leq\sup_{t\in A_{z}^{c}\cap K}\big(((\rho\circ\alpha_{z,\ell})(t))\,(w_{z,\ell}(t))\big)\|\mu\|$
+and
+$\Big\|\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\big((\rho\circ\alpha_{z,s}^{-1})(t)\big)\,\big((w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\big)\|\nu_{s}\|$
+for all $z\in F$ and $s,\ell\in\{1,\dots,N\}$. For each $z\in F$, set
+$\psi_{z}=\widetilde{\mu}_{z}+\frac{\sqrt{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}{\sqrt{\displaystyle\sum_{l=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\Big)}}\;\sum_{s=1}^{N}S^{*}_{z,s}(\widetilde{\nu}_{z,s}).$
+By the choice of $\varepsilon$, the triangle inequality we can deduce that $\psi_{z}\in O$ and
+$\quad\sqrt{\frac{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})(t)\Big)}{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}\;T^{*}_{z,s}(\psi_{z})\in V_{s},$
+for all $z\in F$ and $s\in\{1,\dots,N\}$. This proves the implication $(2)\Rightarrow(1)$. ∎
+Acknowledgement: I am grateful to Professor S. M. Tabatabaie for suggesting linear dynamics on solid spaces as the topic for my research and for introducing to me the relevant literature.